### v1

# Attention In Geometry: An Adaptive Density Field (ADF) with FAISS-Accelerated Spatial Attention

Abstract
Spatial computation in geographic and urban systems increasingly demands scalable mechanisms for aggregating local influence under metric constraints. Classical approaches such as kernel density estimation [silverman1986density] and spatial interpolation either rely on global summation or treat approximation solely as an implementation concern, limiting their interpretability and scalability in large-scale settings. This work introduces Adaptive (Attention) Density Fields (ADF), a geometric attention framework that formulates spatial aggregation as a query-conditioned, metric-induced attention operator in continuous space. Given a set of spatial points with associated scalar scores, ADF constructs a local adaptive Gaussian mixture centered at the query’s nearest neighbors, where kernel bandwidths are modulated by point-specific scores. Approximate nearest-neighbor search, implemented via FAISS [johnson2017faiss] inverted file indices, is treated as an intrinsic component of the attention mechanism, enabling scalable execution while preserving locality. ADF bridges concepts from adaptive kernel methods and attention mechanisms by reinterpreting spatial influence as geometry-preserving attention, grounded directly in physical distance rather than learnt latent projections. The framework is formulation-level rather than algorithm-specific, allowing flexible kernel choices, score-to-bandwidth mappings, and approximation parameters. As such, ADF provides a unifying perspective on spatial influence modeling that emphasizes structure, scalability, and geometric interpretability, with relevance to geographic information systems, urban analysis, and spatial machine learning.
This manuscript is a preprint and has not undergone peer review.
1. Introduction
Large-scale spatial analysis increasingly involves aggregating millions of spatially distributed points of interest (POIs), each associated with heterogeneous influence scores. A common goal is to evaluate a continuous influence field $F(x)$ that reflects the accumulated effect of nearby points under physical distance constraints. However, naive kernel-based summation over all points is computationally infeasible at scale, and classical spatial interpolation methods often treat approximation purely as an implementation concern rather than a structural component of the model.
This work proposes a geometric attention framework for scalable continuous spatial aggregation. Instead of viewing spatial influence as a global density estimation problem, we formulate it as a query-conditioned, metric-induced attention operator in continuous space. Locality, sparsity, and approximation are treated as intrinsic to the formulation, rather than as post hoc accelerations.
This perspective is related in spirit to kernelized and geometry-aware attention mechanisms, but differs in that attention weights are induced directly by physical distance rather than learned projections. [bronstein2021geometric] [choromanski2020rethinking]
Given a set of POIs with associated scalar scores, the framework proceeds as follows:
  • Map POIs from geodetic $(\lambda,\varphi,h)$ to ECEF $\mathbf{x}\in\mathbb{R}^{3}$;
  • Associate each POI with a scalar score $s_{i}$ representing application-dependent influence;
  • Construct an inverted file (IVF) index to accelerate approximate $k$-nearest-neighbor (k-NN) search in 3D;
  • For any query location $\mathbf{x}$:
    • retrieve its $k$ nearest neighbors,
    • define a Gaussian kernel centered at each neighbor, with bandwidth modulated by its score,
    • aggregate the weighted kernel contributions.
The resulting field $F(\mathbf{x})$ constitutes an adaptive Gaussian density/intensity field (ADF): a finite, query-dependent Gaussian mixture that enables scalable evaluation of continuous spatial influence over millions of points with sublinear query complexity.
Importantly, this work focuses on the formulation of geometric attention for spatial aggregation rather than any single algorithmic instantiation. The choice of kernel family, score-to-bandwidth mapping, normalization, and approximation parameters are treated as design degrees of freedom. The emphasis is on the structural decomposition of spatial attention—query-conditioned neighbor selection, metric-grounded weighting, and efficient execution—providing a unifying perspective on adaptive kernel methods, spatial indexing, and attention mechanisms in continuous space.
2. Framework
Figure 1: An overview of the proposed framework
2.1. Overview
The Adaptive Density Field (ADF) is a geometric attention operator defined over continuous three-dimentional space. Given a set of spatial points of interest (POIs) with associated scalar scores, ADF constructs a continuous influence field by aggregating local contributions from nearby points.
For each query location, ADF performs a query-conditioned neighbor selection using an approximate k-NN search, followed by a score-modulated kernel weighting in metric space. The aggregation is explicitly localized, sparse, and adaptive, with kernel bandwidths determined by POI scores rather than fixed global parameters.
To ensure scalability to larger datasets, ADF integrates approximate nearest-neighbor search via FAISS inverted file indices, reducing per query complexity from linear to sublinear in the number of POIs. The resulting operator produces a continuous scalar field that can be evaluated at arbitrary spatial locations and serves as a geometry-preserving attention mechanism grounded in physical distance.
2.2. Data Preparation
Suggested data structures: positions and scores
In the experiments, POIs are derived from a physics-informed trajectory analysis pipeline that detects anomalous motion based on time-normalized Mahalanobis [mahalanobis1936distance] residuals. Each POI is associated with a spatial location and a scalar score reflecting deviation magnitude. The details of this pipeline are not central to the ADF formulation and are therefore omitted; ADF operates on arbitrary spatial point sets with associated scores, defined as follows:
  • $\mathbf{x}_{i}\in\mathbb{R}^{3}$: ECEF position of POI $i$
  • $s_{i}$: score (weight) of POI $i$ Formally:
    $\{\mathbf{x}_{i},s_{i}\}_{i=1}^{n}$
    with:
    $\mathbf{x}_{i}\in\mathbb{R}^{3},\quad s_{i}\in\mathbb{R}.$
    The score $s_{i}$ represents an application-dependent measure of influence, saliency, or importance, and its semantic interpretation is not constrained by the ADF formulation.
2.3. Methodolody
With the POI data, we can perform the ADF construction with the following method:
2.3.1. FAISS IVF Index: Accelerating Neighbor Search
For each POI data we imported, we assign the IVF index [jegou2011product] , mathematically:
Training:
  • FAISS runs k-means with $n_{\text{list}}=4096$ clusters in $\mathbb{R}^{3}$.
  • It learns centroids $\{\mathbf{c}_{j}\}_{j=1}^{4096}$.
    $\min_{\{\mathbf{c}_{j}\}}\sum_{i=1}^{n}\min_{1\leq j\leq 4096}\|\mathbf{x}_{i}-\mathbf{c}_{j}\|^{2}.$
Assignment: Each vector $\mathbf{x}_{i}$ is assigned to its nearest centroid:
$\ell(i)=\arg\min_{1\leq j\leq 4096}\|\mathbf{x}_{i}-\mathbf{c}_{j}\|^{2}.$
and it is stored in inverted list $L_{\ell(i)}$.
Search: For a query $\mathbf{x}$, FAISS finds the 16 nearest centroids:
$j_{1},\dots,j_{16}=\text{indices of the 16 closest }\mathbf{c}_{j}\text{ to }\mathbf{x}.$
and Searches only in the union of lists:
$L_{j_{1}}\cup\dots\cup L_{j_{16}}$
2.3.2. ADF model: an adaptive gaussian mixture
Neighbor search
For a query point $\mathbf{x}\in\mathbb{R}^{3}$ FAISS returns indices:
$\{i_{1},\dots,i_{k}\}=\text{k-NN}_{\text{IVF}}(\mathbf{x})$
in approximate nearest‑neighbor sense. Then set:
  • Neighbor positions:
    $\mathbf{x}_{i_{j}}\in\mathbb{R}^{3}$
  • Neighbor scores:
    $s_{i_{j}}$
Differences
Mathematically, for each neighbor $j=1\dots k$:
$\mathbf{d}_{j}=\mathbf{x}_{i_{j}}-\mathbf{x}=\begin{bmatrix}x_{i_{j},1}-x_{1}\\ x_{i_{j},2}-x_{2}\\ x_{i_{j},3}-x_{3}\end{bmatrix}\in\mathbb{R}^{3}.$
Adaptive Bandwidth
For each neighbor $j$:
$\sigma_{j}=\frac{\sigma_{0}}{s_{i_{j}}+10^{-6}}.$
The $\epsilon=10^{-6}$ term ensures numerical stability and avoids singular bandwidths. Therefore:
  • High score $s_{i_{j}}\implies$ smaller $\sigma_{j}\implies$ more peaked kernel.
  • Low score $s_{i_{j}}\implies$ larger $\sigma_{j}\implies$ broader kernel. This is why it’s adaptive: kernel width depends on the score. Then for each $j$ we have:
    $\text{inv\_var}_{j}=\frac{1}{\sigma_{j}^{2}}.$
NOTE: The specific functional form relating scores to bandwidths (Eq. 5) is a design choice, not a defining property of ADF. ADF is defined by (i) query-conditioned neighbor selection, (ii) metric-induced kernel weighting, and (iii) scalable approximate execution. Any monotonic or learned score-to-bandwidth mapping may be substituted without altering the underlying framework.This work focuses on the general framework of adaptive, score-modulated kernels; alternative monotonic mappings or learned parameterizations can be substituted without altering the structure of ADF.
Quadratic Form (Mahalanobis Distance Squared)
From equation 4 (Eq. 4) we can derive the componentwise square:
$(\mathbf{d}_{j})^{2}=\begin{bmatrix}(x_{i_{j},1}-x_{1})^{2}\\ (x_{i_{j},2}-x_{2})^{2}\\ (x_{i_{j},3}-x_{3})^{2}\end{bmatrix}$
Additionally, we have an array with shape $(k,1)$, broadcasts $\text{inv\_var}_{j}$ across the 3 dimensions. So for each neighbor $j$:
$\text{quad}_{j}=\sum_{d=1}^{3}(x_{i_{j},d}-x_{d})^{2}\cdot\frac{1}{\sigma_{j}^{2}}=\frac{\|\mathbf{x}_{i_{j}}-\mathbf{x}\|^{2}}{\sigma_{j}^{2}}.$
This is exactly the Mahalanobis distance squared for an isotropic Gaussian with variance $\sigma_{j}^{2}$:
$\text{quad}_{j}=(\mathbf{x}-\mathbf{x}_{i_{j}})^{T}\Sigma_{j}^{-1}(\mathbf{x}-\mathbf{x}_{i_{j}}),$
where
$\Sigma_{j}=\sigma_{j}^{2}I_{3},\quad\Sigma_{j}^{-1}=\frac{1}{\sigma_{j}^{2}}I_{3}.$
Gaussian Kernel and ADF Value
For each neighbor $j$, the kernel contribution is:
$K_{j}(\mathbf{x})=\exp\left(-\frac{1}{2}\text{quad}_{j}\right)=\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right).$
In this work, we assume isotropic kernels with $\Sigma_{j}=\sigma_{j}^{2}I_{3}$. And anisotropic kernels are discussed as a future extension. Then the ADF value at $\mathbf{x}$ is:
$F(\mathbf{x})=\sum_{j=1}^{k}s_{i_{j}}\,K_{j}(\mathbf{x})=\sum_{j=1}^{k}s_{i_{j}}\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right).$
Therefore, we claim that ADF is a finite Gaussian mixture [bishop2006pattern] centered at the nearest neighbors, with:
  • centers $\mathbf{x}_{i_{j}}$
  • weights $s_{i_{j}}$
  • bandwidths $\sigma_{j}=\sigma_{0}/(s_{i_{j}}+10^{-6})$.
Assume (i) $s_{i}\geq 0$, or $s_{i}$ is transformed via $s_{i}\leftarrow\text{softplus}(s_{i})$, (ii) $\sigma_{0}$ or the score-to-bandwidth mapping can be learned via gradient descent, (iii) $\sigma_{0}$ has units of meters and controls the physical spatial influence scale.
This can be interpreted as an adaptive kernel density / intensity field in 3D ECEF space.
The contribution of this work lies in the formulation of ADF as a geometric attention mechanism with scalable execution. Kernel choices, bandwidth parameterizations, and score transformations are intentionally left flexible to emphasize the generality of the framework.
2.4. Computational complexity analysis
2.4.1. Time complexity
Without FAISS:
$O(n)\text{ per query}$
With FAISS IVF:
$O(n_{\text{probe}}\cdot\frac{n}{n_{\text{list}}}+k)$
2.4.2. Memory Complexity
Memory complexity is
$O(n)$
dominated by the inverted lists.
2.5. Evaluation
2.5.1. Approximation Error Acknowledgement
Since FAISS IVF performs approximate nearest-neighbor search, ADF(x) is an approximation of the full kernel sum. However, locality of the Gaussian kernel ensures distant errors contribute negligibly. In practice, approximation quality is controlled by $n_{\text{probe}}$ and $k$.
2.5.2. Evaluating ADF at all POIs
For each POI position $\mathbf{x}_{i}$:
$\text{ADF}_{i}=F(\mathbf{x}_{i})=\sum_{j=1}^{k}s_{i_{j}}\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}_{i}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right),$
where $\{i_{1},\dots,i_{k}\}$ are the $k$-nearest neighbors of $\mathbf{x}_{i}$ according to the IVF‑accelerated FAISS search. So the final result is a scalar ADF field defined on all your POIs:
$\mathbf{x}_{i}\mapsto\text{ADF}_{i}.$
2.6. Comparison
2.6.1. Relation to Kernel Density Estimation
Classical KDE [silverman1986density] evaluates a global sum over all points using fixed or locally adaptive bandwidths. In contrast, ADF is explicitly defined as a local, query-dependent aggregation mechanism with bounded support induced by k-NN search. Moreover, ADF is interpreted as a spatial attention operator rather than a probability density estimator, and is designed for scalable execution via approximate nearest-neighbor indices.
ANN-accelerated KDE methods primarily focus on speeding up density estimation. ADF differs in that approximation is not merely an implementation detail, but an explicit component of the attention mechanism, where k-NN selection corresponds to top-k attention pruning.
2.6.2 Relation to the Attention Mechanism
[vaswani2017attention]
While attention mechanisms are typically defined over discrete tokens in a learned latent space, ADF operates directly in continuous metric space. Similarity is induced by physical distance rather than learned projections, and attention weights arise from energy-based kernels instead of normalized dot products. As a result, ADF should be viewed not as a rebranding of Transformer attention, but as a geometric generalization of attention to continuous spatial domains.
Table 1
Here’s the direct correspondence:
ADF Component Attention Component Meaning
Query point $x$ Query vector $Q$ “Where should we look?”
POI positions $x_{i}$ Key vectors $K_{i}$ “What do we compare against?”
POI scores $s_{i}$ Value vectors $V_{i}$ “What do we aggregate?”
Gaussian kernel (Unormalized) attention weights “How strongly do we attend to each neighbor?”
FAISS k‑NN Top‑k attention pruning “Only attend to the nearest keys.”
This correspondence is structural rather than metaphorical:
ADF satisfies the abstract definition of attention as a metric-induced aggregation mechanism over continuous space.
The attention analogy is used as a structural lens rather than a claim of architectural novelty.
Unlike Transformer’s attention, which operates in a learned latent space, ADF attention is geometry-preserving and grounded in physical distance. No projection matrices are required; relevance emerges directly from the spatial metric. Moreover, ADF attention is metric-induced and geometry-preserving, requiring no learned projection matrices.
2.6.3 Why This Matters
The suggested ADF is a geometric attention:
  • Query = location where you want influence
  • Keys = POI positions
  • Values = POI scores
  • Kernel = attention weights
  • FAISS = top‑k attention pruning
This built a spatial attention mechanism.
While ADF shares mathematical components with adaptive KDE and ANN-accelerated density estimation, its contribution is not a new kernel estimator per se. Instead, ADF explicitly formulates these components as a geometric attention operator, where approximation, sparsification, and metric structure are intrinsic to the definition rather than implementation details.
Importantly, the attention analogy is used as a structural lens rather than a claim of architectural novelty. ADF does not seek to replace or subsume Transformer-style attention, but to extend the abstract notion of query-conditioned weighted aggregation to continuous metric spaces.
Overall, ADF should be understood as a framework-level operator rather than a fixed algorithm, with emphasis on structure, scalability, and geometric grounding.
3. Implementation and results
Given a large set of aircraft trajectories, local motion irregularities are first detected as Points of interest (POIs). These POIs are sparse, noisy, and heterogeneous. We aim to aggregate them into a continuous spatial field to reveal Zones of Influence (ZOIs): regions where systematic maneuvering, airspace structure, or control-induced behavior recurrently concentrates.
In many spatial decision problems (e.g., flight routing, navigation, risk exposure), the agent does not respond to absolute field intensity, but to relative dominance along its trajectory. Zones of Influence (ZOIs) are therefore introduced as trajectory-conditioned dominance regions, representing spatial areas where the accumulated influence field locally dominates competing influences relative to the agent’s path.
3.1. Study area and data description
We used a nationwide flight trajectory dataset over mainland China for two multi-periods: 2024-11-10 to 2024-11-22 and 2024-12-05 to 2024-12-15 to extract the POIs. There are a total of 1.8 millions of POIs extracted from the dataset using the motion-based procedure described in Section 3.2.
For the case study on trajectory-conditioned ZOIs, we selected a single day and airport: the Chengdu Shuangliu airport (CTU) on 2024-12-16. All trajectories intersecting the Chengdu region on that day were used to construct a local example of the Adaptive Density Field (ADF) and to illustrate ZOI extraction along individual flights.
The ADF framework is not restricted to flight trajectories and can be instantiated in any domain where localized motion irregularities or events are observed. Likewise, the POI definition is not universal; the construction adopted here is a concrete example tailored to aircraft motion.
3.2. POI construction
To collect the POI data, we applied a mixture of motion-based POI detection methods, detailed in the Appendix. However, POIs are derived from trajectory segments where a physics‑based motion model fails to predict future positions accurately. The processing pipeline operates at the individual flight level. For each flight, raw WGS‑84 positions (longitude, latitude, altitude) are first converted to a local East–North–Up (ENU) coordinate system. The trajectory is then represented as a dictionary keyed by flight identifier, with three aligned time series:
  • coords: longitude, latitude, altitude
  • vel: velocity components (velocity on each axis)
  • dt: timestamp interval
To estimate future aircraft positions, I applied a physics-based interpolation model that blends two motion predictors:
  1. 1.
    Constant-Acceleration (CA) model: reliable for nearly straight trajectories
  2. 2.
    Cubic Hermite Spline interpolation: smooth and accurate for curved motion
the blending weight is determined by the local curvature of the trajectory [lee2007trajectory]: Low curvature indicates motion is nearly straight, hence CA weights more, and vice versa. This adaptive combination produces a more stable and realistic prediction thanm using either method alone.
To evaluate the quality of the predicted positions, I compute a time‑normalized Mahalanobis loss (This metric captures both the magnitude and directional structure of prediction errors.) for each flight. This metric captures not only the magnitude of prediction errors but also their directional structure, covariance, and temporal spacing. After computing the time‑normalized Mahalanobis loss for each flight, the next step is to identify Points of Interest (POIs)—locations where the prediction error is unusually high [chandola2009anomaly]. These points often correspond to sharp maneuvers, abnormal motion, or sensor irregularities, and they serve as valuable markers for downstream analysis.
However, in this case, a POI does not necessarily correspond to an actual infrastructure feature; it simply marks a point where the motion deviates significantly from our prediction. It can be regard as an example tailored POI definition under certain circumstances.
Later we normalize the loss scores and filter the POI with a threshold and exported the data into csv format.
3.3. ADF construction
We instantiated the ADF framework on the nationwide POI dataset by specifying the coordinate system, kernel parameterizations, and neighbor search configurations. All POI locations were converted from WGS84 geodetic coordinates to Earth-Centered, Earth-Fixed (ECEF) Cartesian coordinates to ensure metric consistency. POI scores were treated as non-negative scalar probability attention values and were used directly without additionaly normalization.
For the scalable neighbor retrieval, a FAISS inverted file (IVF) index was introduced and constructed over the ECEF coordinates. The index was trained with 4096 clusters ($2^{12}$) clusters, and queries probed the 16 nearest clusters. For each query point the top $k=100$ approximate nearest neighbors were retrieved. The global spatial scale parameter was fixed to $\sigma_{0}=500$ meters for all experiments. ADF values were evaluated using the formulation defined in Section 2, without further modification.
Unless otherwise stated, these parameter choices were held constant throughout the case study.
3.4. Field evaluation
To analyze the ADF along individual trajectories in Chengdu, we:
  1. 1.
    Convert the coordinates: transform all points from geodetic to ECEF.
  2. 2.
    Evaluate ADF: compute $ADF(x_{t})$ at each trajectory point $x_{t}$ using the FAISS-accelerated framework.
  3. 3.
    Re-attach geodetic coordinates: join ADF values back to $(lon,lat,alt)$ for downstream analysis.
This allows us to examine how aircraft trajectories intersect regions of concentrated POIs.
3.5. ZOI extraction
Given the ADF evaluated along a trajectory, we extract Zones of Influence (ZOIs) as segments where the field intensity dominates a baseline [zheng2015trajectory] defined by the flight’s own experience. Let $\{F_{t}\}$ denote the sequence of ADF values along a single trajectory. We define a robust baseline level as the trajectory‑wise median [rousseeuw1993alternatives]:
$F_{base}={median}_{t}(F_{t})$
A point $t$ is labeled as belonging to a ZOI if
$F_{t}\geq\alpha F_{base}$
where $\alpha>1$ is a dominance factor. In the Chengdu case study, we set $\alpha=1.3$, chosen empirically to balance sensitivity and specificity in detecting high-density segments along individual trajectories. The resulting binary mask partitions the trajectory into ZOI and non‑ZOI segments. Trajectories primarily in low-density regions may still exhibit short ZOIs, while trajectories in busy airspace require higher ADF values to exceed the baseline.
Figure 2: Trajectory ADF analysis and ZOI extraction
Visualization
To illustrate the ADF and ZOI extraction:
  • Trajectories are rendered as white lines, tracing the actual flight paths through the airspace.
  • Red dots mark the extracted Zones of Influence (ZOIs)—trajectory points where the ADF value exceeds the dominance threshold relative to the flight’s own baseline.
  • The remaining trajectory points are colored by their Adaptive Density Field (ADF) values (within range 0~30), using a continuous colormap from blue (low intensity) to red (high intensity), with values ranging from 0 to 80.
  • Layered maps allow toggling between POIs, ADF, ZOIs, and trajectory paths, enabling flexible inspection of each component.
These visualizations demonstrate the relationship between the continuous ADF field and trajectory-conditioned ZOIs. High-ADF segments generally coincide with operationally significant areas such as approach corridors, holding patterns, and vectoring zones in the Chengdu terminal. ZOI points, highlighted in red, mark locations where the trajectory locally exceeds its baseline field intensity; some of these points overlap dense ADF regions, representing recurrent maneuvering, while others appear in otherwise sparse areas, potentially reflecting transient anomalies, unique pilot actions, or data artifacts. By layering ADF intensity, ZOI labels, and raw POIs, the visualization distinguishes between structurally recurrent behavior and isolated events, offering a nuanced perspective on spatial influence and motion irregularities.
4. Discussion
The proposed Adaptive Density Field (ADF) framework provides a flexible, scalable mechanism for aggregating sparse, heterogeneous points of interest (POIs) into a continuous spatial field. By coupling trajectory-conditioned analysis with ZOI extraction as a downstream instantiation of ADF, this methodology captures both globally recurrent patterns and locally significant deviations.
4.1. Future application and extensions
  1. 1.
    Operational Airspace Management: The ADF-ZOI pipeline we presented in the example implementation section could assist air traffic controllers in identifying target regions depending on the specific POI definition, such as regions of frequent maneuvering or congestion, enabling data-driven decisions for route planning, holding pattern optimization, and as well as safety monitoring.
  2. 2.
    Predictive Trajectory Analysis: Beyond post-hoc evaluation, the framework can be integrated with predictive models for trajectory planning, anomaly detection, or risk exposure assessment. By defining the POIs, it is possible to derive trajectory-conditioned ZOIs to provide adaptive thresholds for identifying unusual events relative to expected motion patterns.
  3. 3.
    Cross-Domain Applicability: While this study focuses on aircraft trajectories due to the initial problem setting, the methodolody is domain-agnostic. Any scenario involving spatiotemporal events, as long as the spatial characteristics are stable and fixed (in other words, not dynamic), such as maritime traffic, pedestrian movement, vehicle flows, or wildlife tracking – we expect ADF-based spatial attention is able to identify zones of concentrated activities.
  4. 4.
    Integration with Semantic Context: Future work could embed semantic labels, such as airspace type, weather conditions, or operational procedures, into the ADF formulation, yielding geo-semantic density fields. This way, interestingly, aligns even more closely with the application of attention mechanism in spatiotemporal analysis.
  5. 5.
    Real-Time Applications: With optimizations to indexing, neighbor search, or GPU-accelerated computation, the ADF framework could support near-real-time monitoring of live trajectories in drones, robotics, and other autonomous systems, allowing dynamic ZOI identification and interactive visualization for operational decision-making.
4.2. Methodological Reflections
  • The trajectory-conditioned approach ensures that high-density areas are interpreted relative to the agent’s path, avoiding misleading conclusions based solely on absolute field intensity.
  • Parameter choices, such as kernel bandwidth and dominance factor, influence sensitivity and specificity; future studies could explore adaptive or data-driven tuning to optimize ZOI detection across heterogeneous datasets.
  • Visualization remains crucial for interpretation: layered maps combining POIs, ADF intensity, and ZOI labels provide immediate insight into both recurrent behavior and outlier events, but higher-dimensional visualizations or interactive dashboards could further enhance understanding.
  • This study represents a conceptual framework, and some details may therefore be approximate, incomplete, or provisional.
5. Conclusion
This work introduces Adaptive Density Fields (ADF) as a formulation-level geometric attention operator for scalable aggregation of sparse spatial events in continuous metric space. By defining spatial influence through query-conditioned neighbor selection, score-modulated kernel weighting, and approximate nearest-neighbor execution, ADF reframes adaptive kernel aggregation as an intrinsic form of geometry-preserving attention rather than a purely statistical estimator. The primary contribution lies not in a specific kernel choice or application, but in the structural decomposition of spatial aggregation: locality induced by k-NN attention, metric-grounded weighting, and approximation treated as a defining component of the operator. This perspective unifies ideas from adaptive kernel methods, spatial indexing, and attention mechanisms under a common geometric framework that scales to millions of points while remaining interpretable. To demonstrate the utility of this formulation, we presented an example instantiation in the context of aircraft trajectory analysis. Motion-derived Points of Interest were aggregated into a continuous ADF, and trajectory-conditioned Zones of Influence (ZOIs) were extracted using a relative dominance criterion along individual paths. The Chengdu case study illustrates how the resulting fields reveal recurrent airspace structures while distinguishing localized, trajectory-specific deviations. Beyond aviation, the ADF framework is domain-agnostic and applicable to a wide range of spatial and spatiotemporal settings, including urban mobility, maritime traffic, robotics, and environmental monitoring. Future work will explore semantic augmentation of the field, adaptive parameter learning, anisotropic kernels, and real-time deployment, further extending the applicability of geometric attention in continuous spatial systems. Overall, ADF provides a robust and extensible foundation for scalable, interpretable spatial attention, bridging geometric structure and efficient computation in complex spatial environments.
Appendix: Physics-informed Trajectory POI Detection Pipeline
The appendix provides implementation-level details and mathematical formulations supporting the methods described in the main text. Citations follow the same reference list as the main body.
1 Preprocessing the Flight Data
1.1 Coordinate Conversion [hofmann2012global]: WGS84 Geodetic to ECEF
Given:
  • latitude $\varphi$ (rad)
  • longitude $\lambda$ (rad)
  • ellipsoidal height $h$ (m)
  • WGS84 parameters:
    • semi‑major axis $a=6378137.0$
    • flattening rate $f=\frac{1}{298.257223563}$
    • first eccentricity squared $e^{2}=6.69437999014\times 10^{-3}$
First compute the prime vertical radius of curvature:
$N(\varphi)=\frac{a}{\sqrt{1-e^{2}\sin^{2}\varphi}}$
Then ECEF coordinates $(x,y,z)$:
$x$ $=\left(N(\varphi)+h\right)\cos\varphi\cos\lambda$
$y$ $=\left(N(\varphi)+h\right)\cos\varphi\sin\lambda$
$z$ $=\left(N(\varphi)(1-e^{2})+h\right)\sin\varphi$
Hence we get the ENU coordinates.
1.2 Coordinate Conversion: ECEF to ENU Conversion
Pick a reference point (the origin of the local ENU frame in this case) with geodetic coordinates $(\varphi_{0},\lambda_{0},h_{0})$, and compute its ECEF coordinates $(x_{0},y_{0},z_{0})$ using the same equations as above.
For any point with ECEF $(x,y,z)$, define the difference vector:
$\begin{bmatrix}\Delta x\\ \Delta y\\ \Delta z\end{bmatrix}=\begin{bmatrix}x-x_{0}\\ y-y_{0}\\ z-z_{0}\end{bmatrix}$
And given the Rotation matrix and ENU coordinate at reference $(\varphi_{0},\lambda_{0},h_{0})$:
$\mathbf{R}=\begin{bmatrix}\sin\varphi_{0}&\cos\varphi_{0}&0\\ \cos\varphi_{0}\cdot\sin\lambda_{0}&-\sin\varphi_{0}\cdot\sin\lambda_{0}&\cos\lambda_{0}\\ \cos\varphi_{0}\cdot\cos\lambda_{0}&\sin\varphi_{0}\cdot\cos\lambda_{0}&\sin\lambda_{0}\end{bmatrix}$
Therefore we have the calculation:
$\begin{bmatrix}E\\ N\\ U\end{bmatrix}=\begin{bmatrix}\Delta x\\ \Delta y\\ \Delta z\end{bmatrix}\cdot\mathbf{R}$
This is the standard ECEF → ENU transformation used in geodesy and navigation.
1.3 Creating a Dictionary
To organize per‑flight data extracted from each GeoJSON file, we build a dictionary where each flight ID maps to three lists:
  • coords — longitude, latitude, altitude
  • vel — velocity components
  • dt — timestamps
The basic structure looks like this:
flights = dict({ "coords": [], "vel": [], "dt": [] })
In practice, we use a defaultdict so each new flight_id automatically initializes this structure.
2 Position Prediction
To estimate future aircraft positions, I applied a physics‑based interpolation model that blends two motion predictors:
  1. 1.
    Constant‑Acceleration (CA) model [bar2001estimation] — reliable for nearly straight trajectories
  2. 2.
    Cubic Hermite Spline interpolation [de1978practical] — smooth and accurate for curved motion
The spline utilizes local velocity vectors as tangents at each waypoint, providing a geometrically consistent path [virtanen2020scipy] that complements the CA model’s acceleration-based predictions."
The blending weight is determined by the local curvature of the trajectory: - Low curvature → motion is nearly straight → CA dominates
- High curvature → motion bends → spline dominates
This adaptive combination produces a more stable and realistic prediction than using either method alone.
2.1 General Prediction
For each flight:
  • Convert raw coordinates into a consistent Cartesian frame
  • Compute velocity and approximate acceleration
  • Estimate local curvature $k$ using
$k=\frac{\lVert\mathbf{v}\times\mathbf{a}\rVert}{\lVert\mathbf{v}\rVert^{3}}$
  • Compute a flight‑specific smoothing parameter
$\alpha=\frac{\ln 5}{k_{95}}$
where $k_{95}$ is the 95th percentile curvature
- For each timestamp, compute: - Spline prediction using CubicHermiteSpline - Constant‑acceleration prediction - Blend them using $w=e^{-\alpha k}$:
$\hat{p}=w\,p_{\text{CA}}+(1-w)\,p_{\text{spline}}$
This yields a smooth, curvature‑aware prediction for each flight.
2.2 Loss Computation
To evaluate the quality of the predicted positions, I compute a time‑normalized Mahalanobis loss for each flight. This metric captures not only the magnitude of prediction errors but also their directional structure, covariance, and temporal spacing.
The loss is computed in four main steps:
1. Extract Prediction Residuals
For each flight, I compare the predicted positions $\hat{p}_{i}$ with the actual converted coordinates $p_{i}$:
$r_{i}=\hat{p}_{i}-p_{i}$
Only interior points are used [2 : size-2] to avoid boundary artifacts from the spline and acceleration models.
The residuals are then centered:
$\tilde{r}_{i}=r_{i}-\bar{r}$
This removes global bias and ensures the covariance reflects shape rather than offset.
2. Estimate Residual Covariance
The covariance of the centered residuals is computed as:
$\Sigma=\operatorname{Cov}(\tilde{r})+\lambda I$
A small Tikhonov regularization [tikhonov1963solution] term $\lambda=10^{-5}$ stabilizes the inversion of $\Sigma$, especially for nearly collinear trajectories.
The inverse covariance $\Sigma^{-1}$ defines the Mahalanobis geometry of the error space.
3. Compute Mahalanobis Distance
For each residual vector:
$d_{i}=\sqrt{\tilde{r}_{i}^{\top}\Sigma^{-1}\tilde{r}_{i}}$
This distance penalizes errors more strongly along directions where the model is normally precise, and less along directions with naturally higher variance.
4. Normalize by Temporal Spacing
Because timestamps are not uniformly spaced, each error is scaled by a time‑dependent factor:
$t_{i}=\sqrt{\frac{\Delta t_{i}}{\bar{\Delta t}}}$
The final time-relative Mahalanobis loss is:
$L_{i}=\frac{d_{i}}{t_{i}}$
This ensures that predictions made over longer time intervals are not unfairly penalized compared to short‑interval predictions.
3 POI Detection
After computing the time‑normalized Mahalanobis loss for each flight, the next step is to identify Points of Interest (POIs)—locations where the prediction error is unusually high. These points often correspond to sharp maneuvers, abnormal motion, or sensor irregularities, and they serve as valuable markers for downstream analysis.
However, a POI does not always represent an actual infrastructure feature; it simply marks a point where the motion deviates significantly.
The POI detection pipeline consists of three main stages:
3.1 Normalize the Loss Scores
For each flight, the Mahalanobis losses are rescaled to the interval [0, 1]:
$s_{i}=\frac{L_{i}-\min(L)}{\max(L)-\min(L)+\varepsilon}$
This normalization ensures that POI detection is relative to each flight’s own dynamics, making the method robust to differences in scale, speed, or noise across flights.
3.2 Thresholding
Here, I introduced an element called POI score, which indicates how anomalous each point is relative to the rest of the flight.
A point is flagged as a POI if its normalized score exceeds a fixed threshold:
$s_{i}\geq 0.75$
This threshold captures the upper quartile of anomalous behavior while avoiding excessive false positives.
It can be adjusted depending on the desired sensitivity of the detection process.
3.3 Export POIs to CSV
Each detected POI is stored with:
  • flight ID
  • point index
  • longitude, latitude, altitude
  • POI score
All POIs are aggregated into a Pandas DataFrame and exported as a CSV file, enabling further visualization, inspection, or integration into downstream workflows.

### v2

@@
 Abstract
-Spatial computation in geographic and urban systems increasingly demands scalable mechanisms for aggregating local influence under metric constraints.
-Classical approaches such as kernel density estimation [silverman1986density] and spatial interpolation either rely on global summation or treat approximation solely as an implementation concern, limiting their interpretability and scalability in large-scale settings.
-This work introduces Adaptive (Attention) Density Fields (ADF), a geometric attention framework that formulates spatial aggregation as a query-conditioned, metric-induced attention operator in continuous space.
-Given a set of spatial points with associated scalar scores, ADF constructs a local adaptive Gaussian mixture centered at the query’s nearest neighbors, where kernel bandwidths are modulated by point-specific scores.
-Approximate nearest-neighbor search, implemented via FAISS [johnson2017faiss] inverted file indices, is treated as an intrinsic component of the attention mechanism, enabling scalable execution while preserving locality.
-ADF bridges concepts from adaptive kernel methods and attention mechanisms by reinterpreting spatial influence as geometry-preserving attention, grounded directly in physical distance rather than learnt latent projections.
-The framework is formulation-level rather than algorithm-specific, allowing flexible kernel choices, score-to-bandwidth mappings, and approximation parameters.
-As such, ADF provides a unifying perspective on spatial influence modeling that emphasizes structure, scalability, and geometric interpretability, with relevance to geographic information systems, urban analysis, and spatial machine learning.
-This manuscript is a preprint and has not undergone peer review.
-1. Introduction
-Large-scale spatial analysis increasingly involves aggregating millions of spatially distributed points of interest (POIs), each associated with heterogeneous influence scores.
-A common goal is to evaluate a continuous influence field $F(x)$ that reflects the accumulated effect of nearby points under physical distance constraints.
-However, naive kernel-based summation over all points is computationally infeasible at scale, and classical spatial interpolation methods often treat approximation purely as an implementation concern rather than a structural component of the model.
-This work proposes a geometric attention framework for scalable continuous spatial aggregation.
-Instead of viewing spatial influence as a global density estimation problem, we formulate it as a query-conditioned, metric-induced attention operator in continuous space.
-Locality, sparsity, and approximation are treated as intrinsic to the formulation, rather than as post hoc accelerations.
-This perspective is related in spirit to kernelized and geometry-aware attention mechanisms, but differs in that attention weights are induced directly by physical distance rather than learned projections.
-[bronstein2021geometric] [choromanski2020rethinking]
-Given a set of POIs with associated scalar scores, the framework proceeds as follows:
-  • Map POIs from geodetic $(\lambda,\varphi,h)$ to ECEF $\mathbf{x}\in\mathbb{R}^{3}$;
-  • Associate each POI with a scalar score $s_{i}$ representing application-dependent influence;
-  • Construct an inverted file (IVF) index to accelerate approximate $k$-nearest-neighbor (k-NN) search in 3D;
-  • For any query location $\mathbf{x}$:
-    • retrieve its $k$ nearest neighbors,
-    • define a Gaussian kernel centered at each neighbor, with bandwidth modulated by its score,
-    • aggregate the weighted kernel contributions.
-The resulting field $F(\mathbf{x})$ constitutes an adaptive Gaussian density/intensity field (ADF): a finite, query-dependent Gaussian mixture that enables scalable evaluation of continuous spatial influence over millions of points with sublinear query complexity.
-Importantly, this work focuses on the formulation of geometric attention for spatial aggregation rather than any single algorithmic instantiation.
-The choice of kernel family, score-to-bandwidth mapping, normalization, and approximation parameters are treated as design degrees of freedom.
-The emphasis is on the structural decomposition of spatial attention—query-conditioned neighbor selection, metric-grounded weighting, and efficient execution—providing a unifying perspective on adaptive kernel methods, spatial indexing, and attention mechanisms in continuous space.
-2. Framework
+Spatial computation in geographic and urban systems increasingly requires query-conditioned, local, interpretable aggregation under metric constraints.
+Many classical approaches and spatial interpolations rely on global summation and treat approximation primarily as an implementation concern, which limits their interpretability and scalability in large-scale settings.
+We put forward Adaptive Density Field (ADF) as a geometric attention framework that formulates spatial aggregation as a query-conditioned, metric-induced attention operator in continuous space.
+Given a set of labelled spatial points with associated scalar scores, ADF defines a continuous influence/intensity field over space.
+For a given query location, the field value is obtained via a local adaptive Gaussian kernel mixture centered on the query’s nearest neighbors, where kernel bandwidths are modulated by point-specific scores, such that we can evaluate a scalar value representing local aggregated influence.
+Additionally, approximate nearest-neighbor search is introduced as a complement, which can be implemented with various acceleration algorithms and enables scalable execution while preserving locality.
+The proposed ADF bridges concepts from adaptive kernel methods, classical GIS methods, and the attention mechanisms by reinterpreting spatial influence as geometry-embedded attention, grounded directly in physical distance rather than learnt latent projections.
+The proposed framework is formulation-level rather than algorithm-specific, allowing flexible kernel choices, score-to-bandwidth mappings and approximation parameters.
+Thus, this approach provides a unifying perspective on spatial influence modeling that emphasizes structure, scalability, and geometric interpretability, with relevance to geographic information systems, urban analysis, and spatial machine learning.
+1 Introduction
+Geographic Information Science (GIScience) is a multidisciplinary field [14] concerned with developing analytical and predictive methods for specific tasks involved with geographic data.
+While recent advances in computer science have been dominated by the thriving of large language models [6], people often ignore that a large-scale spatial systems also require a query-conditioned, interpretable aggregation method, like the ones proposed for large language models, that respects metric locality while remaining computationally feasible.
+Such requirements arise naturally in a range of GIScience scenarios, including on-demand accessibility estimation in urban environments, real-time assessment of spatial influence around moving objects (e.g., vehicles, drones, or trajectories), point-of-interest impact modeling, adaptive exposure analysis in environmental monitoring, and interactive spatial querying in large-scale geographic databases [33, 21, 23, 29].
+In these settings, spatial influence must be evaluated locally at arbitrary query locations, often under strict latency and memory constraints, which makes global or batch-based aggregation methods impractical [28].
+Classical adaptive KDE [30] and k-NN density estimator [10] provide a statistical foundation for modern variable bandwidth smoothing, but they are typically global or batch estimators and do not treat approximation as part of the operator definition [4].
+Currently, on one hand several recent applied pipelines have attempted to combine ANN acceleration with KDE for large datasets for better scalability [32, 19, 38], and on the other hand, query-driven KDE variants have also been proposed for influence region computation and other domains [2, 15].
+However, these approaches either assume full density estimation or lack explicit score-modulated attention semantics and intrinsic sparsification, which limits their ability to explicitly encode query-conditioned sparsification, score-dependent influence modulation, and operator-level interpretability.
+In particular, approximation is typically treated as an external optimization rather than an intrinsic component of the aggregation operator, and kernel influence is not formulated as an attention-like mechanism that directly reflects spatial locality and point-specific importance.
+As a result, while being robust and well-established, existing methods unavoidably offer limited flexibility in trading off locality, interpretability, and computational efficiency within a unified formulation.
+In this paper, we propose the Adaptive Density Field (ADF), a geometry‑embedded attention operator that makes locality, sparsification, and approximation first‑class design choices.
+In particular, we consider the problem of constructing a continuous influence field $F(x)$ that reflects the accumulated effect of nearby points of interest (POIs) under physical distance constraints.
+This framework formulates spatial aggregation as a query-conditioned, metric-driven weighted sum.
+For any query location $x\in\mathbb{R}^{3}$ (Earth-Centered, Earth-Fixed coordinates), ADF selects a local neighbor set, assigns each neighbor a score-modulated kernel, and aggregates the contributions to produce $F(x)$, using the FAISS inverted file indices [18] for acceleration.
+This perspective is related to kernelized and geometry-aware attention mechanisms, but differs in that attention weights are induced directly by physical distance rather than learned projections [7, 9].
+This work is structured as follows: In Section 2, we introduce the basic methodology and formal definition of the Adaptive Density Field (ADF), positioning the framework as a conceptual contribution to GIScience.
+In Section 3, we present a practical instantiation that combines ADF with ANN indices for POI detection.
+In Section 4 we evaluate the robustness and efficiency of the proposed framework, and demonstrate the selection of parameters.
+In Section 5 we discuss the relationship and differences between the proposed method and several classic and well-known GIS method.
+In Section 6, we discuss the implications of ADF as a unifying operator-level perspective on spatial influence modeling, emphasizing geometric grounding, explicit approximation, and scalability.
+Finally, Section 7 summarizes the main conclusions and outlines future work.
+2 Framework
+2.1 Overview
+The Adaptive Density Field (ADF) is a geometric attention operator designed to unify locality, sparsity, and approximation in continuous spatial aggregation.
+Rather than focusing on implementation or dataset-specific optimizations, the framework formalizes spatial influence as an operator-level construct: given a query location $x\in\mathbb{R}^{3}$ (ECEF coordinates), ADF aggregates contributions from nearby points of interests (POIs) through score-modulated kernels and a metric-induced weighting scheme.
+This section provides a conceptual roadmap of the framework.
+We first define the operator mathematically, discuss the score-to-bandwidth mappings, and approximation mechanisms, and then present the usage of FAISS for efficient neighbor retrieval as an approximate nearest-neighbor (ANN) method.
+Finally, we evaluate the time and memory complexity of the proposed method, highlighting the flexibility of the approach and its suitability for a range of GIScience applications.
+2.2 Data Preparation
+Suggested data structures: positions and scores
+In the experiments, POIs are derived from a physics-informed trajectory analysis pipeline (detailed in the Appendix); each POI is associated with a spatial location and a scalar score reflecting deviation magnitude.
+The details of this pipeline are not central to the ADF formulation and are therefore omitted.
+ADF operates on arbitrary spatial point sets with associated scores, defined as follows:
+(i) $\mathbf{x}_{i}\in\mathbb{R}^{3}$ is the ECEF position of POI $i$
+(ii) $s_{i}$ is the score (weight) of POI $i$.
+Formally, the input is defined as $\{\mathbf{x}_{i},s_{i}\}_{i=1}^{n}$, where $\mathbf{x}_{i}\in\mathbb{R}^{3}$ and $s_{i}\in\mathbb{R}$.
+The score $s_{i}$ represents an application-dependent measure of influence, saliency, or importance; its semantic interpretation is not constrained by the ADF formulation.
 Figure 1: An overview of the proposed framework
-2.1. Overview
-The Adaptive Density Field (ADF) is a geometric attention operator defined over continuous three-dimentional space.
-Given a set of spatial points of interest (POIs) with associated scalar scores, ADF constructs a continuous influence field by aggregating local contributions from nearby points.
-For each query location, ADF performs a query-conditioned neighbor selection using an approximate k-NN search, followed by a score-modulated kernel weighting in metric space.
-The aggregation is explicitly localized, sparse, and adaptive, with kernel bandwidths determined by POI scores rather than fixed global parameters.
-To ensure scalability to larger datasets, ADF integrates approximate nearest-neighbor search via FAISS inverted file indices, reducing per query complexity from linear to sublinear in the number of POIs.
-The resulting operator produces a continuous scalar field that can be evaluated at arbitrary spatial locations and serves as a geometry-preserving attention mechanism grounded in physical distance.
-2.2. Data Preparation
-Suggested data structures: positions and scores
-In the experiments, POIs are derived from a physics-informed trajectory analysis pipeline that detects anomalous motion based on time-normalized Mahalanobis [mahalanobis1936distance] residuals.
-Each POI is associated with a spatial location and a scalar score reflecting deviation magnitude.
-The details of this pipeline are not central to the ADF formulation and are therefore omitted; ADF operates on arbitrary spatial point sets with associated scores, defined as follows:
-  • $\mathbf{x}_{i}\in\mathbb{R}^{3}$: ECEF position of POI $i$
-  • $s_{i}$: score (weight) of POI $i$ Formally:
-    $\{\mathbf{x}_{i},s_{i}\}_{i=1}^{n}$
-    with:
-    $\mathbf{x}_{i}\in\mathbb{R}^{3},\quad s_{i}\in\mathbb{R}.$
-    The score $s_{i}$ represents an application-dependent measure of influence, saliency, or importance, and its semantic interpretation is not constrained by the ADF formulation.
-2.3. Methodolody
-With the POI data, we can perform the ADF construction with the following method:
-2.3.1. FAISS IVF Index: Accelerating Neighbor Search
-For each POI data we imported, we assign the IVF index [jegou2011product] , mathematically:
-Training:
-  • FAISS runs k-means with $n_{\text{list}}=4096$ clusters in $\mathbb{R}^{3}$.
-  • It learns centroids $\{\mathbf{c}_{j}\}_{j=1}^{4096}$.
-    $\min_{\{\mathbf{c}_{j}\}}\sum_{i=1}^{n}\min_{1\leq j\leq 4096}\|\mathbf{x}_{i}-\mathbf{c}_{j}\|^{2}.$
-Assignment: Each vector $\mathbf{x}_{i}$ is assigned to its nearest centroid:
-$\ell(i)=\arg\min_{1\leq j\leq 4096}\|\mathbf{x}_{i}-\mathbf{c}_{j}\|^{2}.$
+2.3 Methodology
+Given the POI data, we construct the ADF as follows:
+2.3.1 FAISS IVF Index: Accelerating Neighbor Search
+For the proposed framework, FAISS is used purely as a computational accelerator to retrieve local neighborhoods efficiently from large POI sets.
+Parameters such as k and nprobe control access to sufficient local context rather than defining the ADF model itself (we will discuss the parameter selection in Section 4).
+In practice, results were stable once neighborhoods exceeded a minimal size, and all experiments used a fixed, conservative configuration.
+For each POI data imported, we assign the IVF index [17], and the procedure for high-dimensional indexing is formalized as follows:
+Training
+(i) FAISS runs k-means with $n_{\text{list}}=4096$ clusters in $\mathbb{R}^{3}$.
+(ii) It learns centroids $\{\mathbf{c}_{j}\}_{j=1}^{4096}$.
+This formalizes the cluster assignment, though standard FAISS documentation may be consulted for full details.
+Assignment
+Each vector $\mathbf{x}_{i}$ is assigned to its nearest centroid:
+$\ell(i)=\arg\min_{1\leq j\leq 4096}\|\mathbf{x}_{i}-\mathbf{c}_{j}\|^{2}.$ (1)
 and it is stored in inverted list $L_{\ell(i)}$.
-Search: For a query $\mathbf{x}$, FAISS finds the 16 nearest centroids:
-$j_{1},\dots,j_{16}=\text{indices of the 16 closest }\mathbf{c}_{j}\text{ to }\mathbf{x}.$
-and Searches only in the union of lists:
-$L_{j_{1}}\cup\dots\cup L_{j_{16}}$
-2.3.2. ADF model: an adaptive gaussian mixture
+Search
+For a query $\mathbf{x}$, FAISS finds the 16 nearest centroids $j_{1},\dots,j_{16}$: indices of the 16 closest $\mathbf{c}_{j}$ to $\mathbf{x}.$ Then searches only in the union of these lists: $L_{j_{1}}\cup\dots\cup L_{j_{16}}$.
+2.3.2 ADF Model: An Adaptive Gaussian Mixture
 Neighbor search
-For a query point $\mathbf{x}\in\mathbb{R}^{3}$ FAISS returns indices:
-$\{i_{1},\dots,i_{k}\}=\text{k-NN}_{\text{IVF}}(\mathbf{x})$
-in approximate nearest‑neighbor sense. Then set:
-  • Neighbor positions:
-    $\mathbf{x}_{i_{j}}\in\mathbb{R}^{3}$
-  • Neighbor scores:
-    $s_{i_{j}}$
+For a query point $\mathbf{x}\in\mathbb{R}^{3}$ FAISS returns indices: $\{i_{1},\dots,i_{k}\}=\text{k-NN}_{\text{IVF}}(\mathbf{x})$ in approximate nearest‑neighbor sense.
+Then set: (i) Neighbor positions: $\mathbf{x}_{i_{j}}\in\mathbb{R}^{3}$; (ii) Neighbor scores: $s_{i_{j}}$
 Differences
-Mathematically, for each neighbor $j=1\dots k$:
-$\mathbf{d}_{j}=\mathbf{x}_{i_{j}}-\mathbf{x}=\begin{bmatrix}x_{i_{j},1}-x_{1}\\ x_{i_{j},2}-x_{2}\\ x_{i_{j},3}-x_{3}\end{bmatrix}\in\mathbb{R}^{3}.$
+Mathematically, for each neighbor $j=1\dots k$, we have:
+$\mathbf{d}_{j}=\mathbf{x}_{i_{j}}-\mathbf{x}\in\mathbb{R}^{3}.$ (2)
 Adaptive bandwidth
-For each neighbor $j$:
-$\sigma_{j}=\frac{\sigma_{0}}{s_{i_{j}}+10^{-6}}.$
+Conceptually, higher score points are supposed to represent stronger local influence, hence narrower kernels; lower scores should spread influence more diffusely. Therefore, for each neighbor $j$:
+$\sigma_{j}=\frac{\sigma_{0}}{s_{i_{j}}+10^{-6}}.$ (3)
 The $\epsilon=10^{-6}$ term ensures numerical stability and avoids singular bandwidths. Therefore:
   • High score $s_{i_{j}}\implies$ smaller $\sigma_{j}\implies$ more peaked kernel.
-  • Low score $s_{i_{j}}\implies$ larger $\sigma_{j}\implies$ broader kernel. This is why it’s adaptive: kernel width depends on the score. Then for each $j$ we have:
-    $\text{inv\_var}_{j}=\frac{1}{\sigma_{j}^{2}}.$
-NOTE: The specific functional form relating scores to bandwidths (Eq.
-5) is a design choice, not a defining property of ADF.
-ADF is defined by (i) query-conditioned neighbor selection, (ii) metric-induced kernel weighting, and (iii) scalable approximate execution.
-Any monotonic or learned score-to-bandwidth mapping may be substituted without altering the underlying framework.This work focuses on the general framework of adaptive, score-modulated kernels; alternative monotonic mappings or learned parameterizations can be substituted without altering the structure of ADF.
-Quadratic Form (Mahalanobis Distance Squared)
-From equation 4 (Eq. 4) we can derive the componentwise square:
-$(\mathbf{d}_{j})^{2}=\begin{bmatrix}(x_{i_{j},1}-x_{1})^{2}\\ (x_{i_{j},2}-x_{2})^{2}\\ (x_{i_{j},3}-x_{3})^{2}\end{bmatrix}$
-Additionally, we have an array with shape $(k,1)$, broadcasts $\text{inv\_var}_{j}$ across the 3 dimensions. So for each neighbor $j$:
-$\text{quad}_{j}=\sum_{d=1}^{3}(x_{i_{j},d}-x_{d})^{2}\cdot\frac{1}{\sigma_{j}^{2}}=\frac{\|\mathbf{x}_{i_{j}}-\mathbf{x}\|^{2}}{\sigma_{j}^{2}}.$
-This is exactly the Mahalanobis distance squared for an isotropic Gaussian with variance $\sigma_{j}^{2}$:
-$\text{quad}_{j}=(\mathbf{x}-\mathbf{x}_{i_{j}})^{T}\Sigma_{j}^{-1}(\mathbf{x}-\mathbf{x}_{i_{j}}),$
+  • Low score $s_{i_{j}}\implies$ larger $\sigma_{j}\implies$ broader kernel. This is why it’s adaptive: kernel width depends on the score.
+Then for each $j$ we have:
+$\text{inv\_var}_{j}=\frac{1}{\sigma_{j}^{2}}.$ (4)
+NOTE: The specific functional form relating scores to bandwidths is a design choice, not a defining property of ADF.
+ADF is defined by (i) query-conditioned neighbor selection; (ii) metric-induced kernel weighting; and (iii) scalable approximate execution.
+While this study employs a specific reciprocal mapping, any monotonic or learned parameterization may be substituted without altering the underlying framework.
+Quadratic form (Mahalanobis distance squared [22])
+From Equation 3 we can derive the componentwise square $\mathbf{d}_{j}^{2}$.
+Additionally, we have an array with shape $(k,1)$, broadcasts $\text{inv\_var}_{j}$ across the 3 dimensions.
+So for each neighbor $j$:
+$\text{quad}_{j}=\sum_{d=1}^{3}(x_{i_{j},d}-x_{d})^{2}\cdot\frac{1}{\sigma_{j}^{2}}=\frac{\|\mathbf{x}_{i_{j}}-\mathbf{x}\|^{2}}{\sigma_{j}^{2}}.$ (5)
+This is a derived Mahalanobis distance squared for an isotropic Gaussian with variance $\sigma_{j}^{2}$:
+$\text{quad}_{j}=(\mathbf{x}-\mathbf{x}_{i_{j}})^{T}\Sigma_{j}^{-1}(\mathbf{x}-\mathbf{x}_{i_{j}}),$ (6)
 where
@@
 For each neighbor $j$, the kernel contribution is:
-$K_{j}(\mathbf{x})=\exp\left(-\frac{1}{2}\text{quad}_{j}\right)=\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right).$
-In this work, we assume isotropic kernels with $\Sigma_{j}=\sigma_{j}^{2}I_{3}$. And anisotropic kernels are discussed as a future extension. Then the ADF value at $\mathbf{x}$ is:
-$F(\mathbf{x})=\sum_{j=1}^{k}s_{i_{j}}\,K_{j}(\mathbf{x})=\sum_{j=1}^{k}s_{i_{j}}\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right).$
-Therefore, we claim that ADF is a finite Gaussian mixture [bishop2006pattern] centered at the nearest neighbors, with:
-  • centers $\mathbf{x}_{i_{j}}$
-  • weights $s_{i_{j}}$
+$K_{j}(\mathbf{x})=\exp\left(-\frac{1}{2}\text{quad}_{j}\right)=\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right).$ (7)
+In this work, we assume isotropic kernels with $\Sigma_{j}=\sigma_{j}^{2}I_{3}$.
+While the ADF framework natively supports anisotropic covariance matrices to account for directional trajectory influence, for simplification, we applied isotropic kernels in this study to establish a controlled baseline for evaluating the scalability and partitioning stability of the ADF.
+This isotropic simplification isolates the effects of score-modulated adaptation and scalable neighbor retrieval, establishing a robust baseline before introducing directional anisotropy in future trajectory-specific extensions.
+Then the ADF value at $\mathbf{x}$ is assumed:
+$F(\mathbf{x})=\sum_{j=1}^{k}s_{i_{j}}\,K_{j}(\mathbf{x})=\sum_{j=1}^{k}s_{i_{j}}\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right).$ (8)
+Note that $F(x)$ is not normalized to integrate to unity and is therefore not a probability density.
+Therefore, we claim that ADF is a finite Gaussian mixture [5].
+Here, the term ‘Gaussian mixture’ refers to a computational representation rather than a probabilistic generative model, centered at the nearest neighbors, with:
+  • centers $\mathbf{x}_{i_{j}}$;
+  • weights $s_{i_{j}}$;
   • bandwidths $\sigma_{j}=\sigma_{0}/(s_{i_{j}}+10^{-6})$.
-Assume (i) $s_{i}\geq 0$, or $s_{i}$ is transformed via $s_{i}\leftarrow\text{softplus}(s_{i})$, (ii) $\sigma_{0}$ or the score-to-bandwidth mapping can be learned via gradient descent, (iii) $\sigma_{0}$ has units of meters and controls the physical spatial influence scale.
-This can be interpreted as an adaptive kernel density / intensity field in 3D ECEF space.
+We assume (i) $s_{i}\geq 0$, or $s_{i}$ is transformed via $s_{i}\leftarrow\text{softplus}(s_{i})$, (ii) $\sigma_{0}$ or the score-to-bandwidth mapping can be learned via gradient descent, (iii) $\sigma_{0}$ has units of meters and controls the physical spatial influence scale.
+This can be interpreted as an adaptive kernel influence / intensity field in 3D ECEF space.
 The contribution of this work lies in the formulation of ADF as a geometric attention mechanism with scalable execution.
 Kernel choices, bandwidth parameterizations, and score transformations are intentionally left flexible to emphasize the generality of the framework.
-2.4. Computational complexity analysis
-2.4.1. Time complexity
-Without FAISS:
-$O(n)\text{ per query}$
-With FAISS IVF:
-$O(n_{\text{probe}}\cdot\frac{n}{n_{\text{list}}}+k)$
-2.4.2. Memory Complexity
-Memory complexity is
-$O(n)$
-dominated by the inverted lists.
-2.5. Evaluation
-2.5.1. Approximation Error Acknowledgement
+In future extensions tailored to directional flow data (e.g., aircraft trajectories), anisotropic covariances can be induced by aligning principal axes with local velocity vectors, fully realizing vector-aware geometric attention absent in traditional VBKDE.
+2.3.3 Evaluating ADF at all POIs
+From Equation 8, for each POI position $\mathbf{x}_{i}$:
+$\text{ADF}_{i}=\sum_{j=1}^{k}s_{i_{j}}\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}_{i}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right),$ (9)
+where $F(x)$ evaluated at POI $x_{i}$ is denoted as $ADF_{i}$, and $\{i_{1},\dots,i_{k}\}$ are the $k$-nearest neighbors of $\mathbf{x}_{i}$ according to the IVF‑accelerated FAISS search.
+So the final result is a scalar ADF field defined on all your POIs:
+$\mathbf{x}_{i}\mapsto\text{ADF}_{i}.$
+2.3.4 Approximation Error Acknowledgement
 Since FAISS IVF performs approximate nearest-neighbor search, ADF(x) is an approximation of the full kernel sum.
@@
 In practice, approximation quality is controlled by $n_{\text{probe}}$ and $k$.
-2.5.2. Evaluating ADF at all POIs
-For each POI position $\mathbf{x}_{i}$:
-$\text{ADF}_{i}=F(\mathbf{x}_{i})=\sum_{j=1}^{k}s_{i_{j}}\exp\left(-\frac{1}{2}\frac{\|\mathbf{x}_{i}-\mathbf{x}_{i_{j}}\|^{2}}{\sigma_{j}^{2}}\right),$
-where $\{i_{1},\dots,i_{k}\}$ are the $k$-nearest neighbors of $\mathbf{x}_{i}$ according to the IVF‑accelerated FAISS search. So the final result is a scalar ADF field defined on all your POIs:
-$\mathbf{x}_{i}\mapsto\text{ADF}_{i}.$
-2.6. Comparison
-2.6.1. Relation to Kernel Density Estimation
-Classical KDE [silverman1986density] evaluates a global sum over all points using fixed or locally adaptive bandwidths.
-In contrast, ADF is explicitly defined as a local, query-dependent aggregation mechanism with bounded support induced by k-NN search.
-Moreover, ADF is interpreted as a spatial attention operator rather than a probability density estimator, and is designed for scalable execution via approximate nearest-neighbor indices.
-ANN-accelerated KDE methods primarily focus on speeding up density estimation.
-ADF differs in that approximation is not merely an implementation detail, but an explicit component of the attention mechanism, where k-NN selection corresponds to top-k attention pruning.
-2.6.2 Relation to the Attention Mechanism
-[vaswani2017attention]
+Detailed numerical results for precision and recall are provided in Sections 4 and 5.
+2.4 Computational Complexity Analysis
+To assess the computational feasibility of the ADF framework for large-scale trajectory analysis, we provide a theoretical analysis of its time and memory complexity.
+Detailed benchmark experiments and comparison are presented in Section 5.2.
+Here, we provide theoretical estimates to illustrate the framework’s scalability potential for global POI datasets.
+2.4.1 Time Complexity
+The standard brute-force approach for calculating influence across $n$ POIs would require $O(n)$ distance computations per query point. In contrast, the FAISS IVF index reduces this to:
+$O(n_{probe}\cdot\frac{n}{n_{list}}+k)$ (10)
+where $n_{probe}$ is the number of centroids searched, $n/n_{list}$ represents the average size of an inverted list, and $k$ is the cost of sorting the final candidates.
+For large-scale datasets where $n_{list}\approx\sqrt{n}$, this effectively transforms the search into sub-linear time, enabling real-time influence field construction for high-frequency trajectory data.
+2.4.2 Memory Complexity
+The memory overhead is dominated by the inverted lists storing the ECEF coordinates. The complexity is:
+$O(n\cdot d)$ (11)
+where $d=3$ for our 3D spatial vectors.
+Given that each POI requires around 12 bytes (3 floats), the framework is theoretically capable of accommodating millions of POIs on commodity hardware with several gigabytes of RAM.
+This demonstrates that the ADF framework balances expressive modeling with computational practicality, making it suitable for global-scale geospatial applications.
+3 Case Study: Flight Trajectory POI Extraction
+The ADF framework is applicable to a wide range of applications, which we will discuss in the Section 6.
+To illustrate the framework’s utility, we evaluate the extraction of the POIs from flight trajectories using the ADF method compared against the supervised kinematic baseline (the suggested grounded truth) introduced in the Appendix.
+In this setup we used a nationwide flight trajectory dataset over mainland China as the reference point set for two multi-day time periods: 2024-11-10 to 2024-11-22 and 2024-12-05 to 2024-12-16 (except the Chengdu Shaungliu Airport on 2024-12-16, which is used for evaluation) to extract the POIs.
+A total of approximately 1.8 million POIs were extracted from the dataset using the motion-based procedure described in Section 3.1.
+Additionally, we selected a single day and airport: the Chengdu Shuangliu International Airport (CTU) on 2024-12-16.
+The mentioned nationwide POI were used to construct a local example of the Adaptive Density Field (ADF) and to illustrate POI extraction along all the trajectories arriving CTU that day.
+3.1 Baseline POI Construction
+To collect the POI data, we applied a physics-based motion residual analysis pipeline as an example of ’manually labelled’ POI baseline.
+POIs are derived from trajectory segments where the motion model fails to predict future positions accurately.
+To estimate future aircraft positions, we applied a physics-based interpolation model that blends two motion predictors:
+  1. 1.
+    Constant-Acceleration (CA) model: reliable for nearly straight trajectories
+  2. 2.
+    Cubic Hermite Spline interpolation: smooth and accurate for curved motion
+the blending weight $w$ is an exponential decay function of the local curvature $\kappa$ [20], calibrated against the 95th percentile of curvatures within each flight: Low curvature indicates motion is nearly straight, hence CA weights more, and vice versa.
+This adaptive combination produces a more stable and realistic prediction than using either method alone.
+To evaluate the quality of the predicted positions, we computed the time‑normalized Mahalanobis loss [22] for each flight.
+This metric captures not only the magnitude of prediction errors but also their directional structure, covariance, and temporal spacing.
+Then we labelled the Points of Interest (POIs)—locations where the prediction error is unusually high [8].
+These points often correspond to sharp maneuvers, abnormal motion, or sensor irregularities, and they serve as valuable markers for downstream analysis.
+Loss scores are normalized to extract POIs, and the threshold in this case study is the $75\%$ percentile.
+However, in this case, the detected POI does not necessarily correspond to an actual infrastructure feature; it simply marks a point where the motion deviates significantly from our prediction.
+It should therefore be understood as an example of task-specific POI definition tailored to aircraft motion.
+Full implementation details and parameter settings are provided in the Appendix.
+3.2 ADF Framework and FAISS Acceleration
+We instantiated the ADF framework on the nationwide POI dataset by specifying the coordinate system, kernel parameterizations, and neighbor search configurations.
+All POI locations were converted from WGS84 [1] geodetic coordinates to Earth-Centered, Earth-Fixed (ECEF) Cartesian coordinates to ensure metric consistency.
+POI scores were treated as non-negative scalar probability attention values and were used directly without additional normalization.
+For the scalable neighbor retrieval, we implemented a FAISS-based Inverted File (IVF) index structure over the ECEF coordinates.
+The spatial domain was partitioned into $4096$ ($2^{12}$) Voronoi cells (clusters) during the training phase using a flat L2 quantizer.
+This partitioning scheme optimizes the search space by narrowing the query range to specific geographic regions rather than entire nationwide dataset.
+During the inference phase, we configured the nprobe parameter to $16$, meaning that for each query point, only the $16$ nearest Voronoi cells, those with centroids closest to the query location were probed.
+This configuration allows for high-fidelity approximate nearest neighbor (ANN) retrieval of the top $k=100$ neighbors while maintaining the computational efficiency required for processing millions of trajectory points.
+Crucially, the global spatial scale parameter was fixed to $\sigma_{0}=500$ meters for all experiments, providing a consistent baseline for the kernel bandwidth.
+The ADF values were then evaluated strictly using the formulation defined in Section 2, ensuring that the adaptive kernel $(\sigma=\sigma_{0}/(scores+1e-6))$ remained mathematically consistent with the established framework without further modification.
+Unless otherwise stated, these parameter choices were held constant throughout the case study.
+3.3 Evaluation and POI Extraction
+To analyze the ADF along individual trajectories in Chengdu region, we:
+  1. 1.
+    Convert the coordinates: Transform all points from WGS84 geodetic coordinated to ECEF Cartesian coordinates.
+  2. 2.
+    Evaluate ADF: Compute the $ADF(x_{t})$ at each trajectory point using the FAISS-accelerated framework with adaptive bandwidth defined in Section 2.
+  3. 3.
+    Re-attach geodetic coordinates: Join the computed ADF values back to the original $(lon,lat,alt,t)$ records for downstream spatial-temporal analysis.
+The extraction of final POIs from the density field follows the principle that influence should be assessed relative to operational context rather than a rigid global threshold [39].
+Because aircraft operate across varying density regimes (e.g., congested terminal areas versus sparse cruise sectors), we employ a trajectory-specific relative threshold.Let $\{F_{t}\}$ denote the sequence of ADF values along a single trajectory.
+A trajectory point at time $t$ is labeled as a POI if:
+$F_{t}\geq P_{75}(\{F_{t}\})$
+where $P_{75}$ represents the 75th percentile of field intensities experienced by that specific flight, this percentile was selected to maintain consistency with the baseline prevalence discussed in Section 3.1.
+As demonstrated in Section 4, utilizing a relative threshold ensures statistical parity between the baseline and ADF results, allowing for a one-to-one comparison of the detection fidelity across the evaluation dataset.
+This adaptive criterion yields a binary mask along each trajectory, effectively normalizing the detection process against the "background noise" of the flight’s environment.
+Consequently, a flight primarily traversing low-density regions can still exhibit localized POIs (e.g., a sudden maneuver in mid-air), while a flight in a persistently congested terminal area must exceed its own higher-than-average baseline to trigger a detection.
+This approach ensures high specificity and prevents the "hallucination" of POIs in areas of high global density that lack local behavioral significance.
+3.4 Comparative Validation: ADF vs. Kinematic Baseline
+The comparative validation demonstrates a high degree of spatial consistency between the ADF framework and the kinematic baseline.
+Notably, as the threshold increases from 150 m to 300 m, the ADF’s matched points rise from 16,657 to 18,074, representing a consistency rate of approximately $80\%$ relative to its own total POI count ($20,769$).
+The observed ratio suggests that the ADF method effectively resolves coarse kinematic anomalies into fine-grained behavioral clusters.
+As the ADF and kinematic baseline rely on distinct mathematical foundations, a strict coordinate-wise overlap is not expected.
+Instead, the validation assesses spatial concordance within a specified distance threshold.
+To ensure a fair comparative analysis, we enforced cardinality matching between the two sets; where the ADF generated multiple candidate points for a single kinematic event, the redundant labels were consolidated to ensure the total population ($N$) remained consistent across both methods.
+This normalization allows us to test the model’s labeling accuracy without the results being skewed by differences in point density.
+Table 1: Comparative Spatial Matching: Baseline vs. ADF (Fine-Scale Sensitivity)
+150 meters 200 meters 300 meters
+Category Base ADF Base ADF Base ADF
+Matched (TT) 16,606 17,225 17,953
+Unique (TF/FT) 10,758 4,163 10,492 3,544 10,056 2,816
+Precision (%) 79.96% 82.94% 86.44%
+Note: Precision is defined as the ratio of matched points to total points ($N_{ADF}=20,769$).
+While Table 1 details the precise point counts for fine-scale thresholds, Figure 2 illustrates the broader convergence of the two methods.
+As the spatial tolerance is relaxed to 500 m, the ADF precision plateaus at above $90\%$, confirming that the vast majority of ADF-extracted POIs are spatially anchored to kinematic anomalies, even if they exhibit higher local precision than the baseline at the 100 m scale.
+Figure 2: ADF Spatial Fidelity Sensitivity Analysis.
+This plot illustrates the performance metrics of the ADF framework, including Precision, Recall, and F1-score, relative to the kinematic baseline across an expanding distance threshold (100 m to 500 m).
+The asymptotic increase in precision beyond 400 m demonstrates that the ADF method anchors to kinematic behavioral anomalies while maintaining high spatial concordance, even as the "soft" matching criteria are relaxed.
+Performance improves smoothly with increasing spatial tolerance, with precision exceeding $80\%$ at 200 m and stabilizing above $90\%$ beyond 400 m, while recall (defined as baseline-covered ADF detections) remains consistently bounded around $60\sim 70\%$, indicating stable spatial localization rather than threshold-dependent artifacts.
+Figure 3: Spatial Distribution of Extracted POIs in Chengdu Region.
+This 2D projection illustrates the geographic distribution of points identified by the ADF framework (cyan) and the kinematic baseline (red).
+While all computations in Section 3, including neighbor retrieval and density field construction, were performed in 3D ECEF Cartesian space to ensure metric integrity, the results are visualized here in 2D geodetic coordinates for regional context.
+The significant overlap in high-traffic corridors and terminal areas near CTU visually confirms the spatial concordance quantified in Table 1.
+It is important to emphasize that while Figure 3 provides a 2D planimetric view of the results for visualization purposes, all underlying experiments and proximity searches were executed in the full 3D ECEF coordinates system.
+This ensures that the vertical dimension of flight trajectories is fully accounted for in the ADF mathematical model, even when projected onto a 2D map.
+4 Ablation Study
+To evaluate the robustness and efficiency of the proposed framework, a series of ablation experiments were conducted.
+This study systematically isolates the key components of the system, computational acceleration via indexing, kernel bandwidth strategies, and hyperparameter sensitivity to quantify their contribution to the overall performance.
+By comparing the full implementation against simplified or non-optimized baselines, we establish the empirical justification for the architectural choices detailed in Section 3.
+4.1 On Score and Bandwidths
+A critical feature of the ADF is its use of adaptive bandwidths ($\sigma$) and POI scores ($s$).
+To verify the necessity of this complexity, the model was compared against a standard fixed-bandwidth Kernel Density Estimation (KDE) where $\sigma$ was held constant (150 and 200) and $s$ was ignored (effectively treating all POIs as equally significant).
+In Table 2 and 3, we demonstrated the comparison between fixed bandwidth (250, 500, 750 meters) and adaptive bandwidth under 150 and 200 meters tolerance threshold (The nprobe for FAISS is set to 16, and k=100).
+While fixed bandwidths in some cases can achieve higher recall by capturing smaller, isolated POI clusters, they all suffer from a slight reduction in precision whose intensity is affected by the bandwidth settings, which make the output unstable and hard to predict.
+The Adaptive ADF utilizes local POI scores to modulate the kernel width, which provides a superior balance between spatial specificity and sensitivity, resulting in the stability of the precision and F1 score.
+Table 2: Comparison table: fixed and adaptive bandwidths (150m threshold)
+Bandwidth $\sigma$ Precision Recall F1
+adaptive $79.96\%$ $60.69\%$ $69.00\%$
+250 $76.98\%$ $61.04\%$ $68.09\%$
+500 $79.77\%$ $60.57\%$ $68.86\%$
+750 $79.84\%$ $60.60\%$ $68.91\%$
+Table 3: Comparison table: fixed and adaptive bandwidths (200m threshold)
+Bandwidth $\sigma$ Precision Recall F1
+adaptive $82.94\%$ $62.15\%$ $71.05\%$
+250 $79.55\%$ $62.44\%$ $69.96\%$
+500 $82.63\%$ $61.97\%$ $70.83\%$
+750 $82.84\%$ $62.06\%$ $70.96\%$
+4.2 On Implementation of FAISS
+To validate the efficiency of the proposed system, a comparative analysis was conducted between the optimized FAISS-based implementation and a traditional brute force approach.
+This experiment aimed to quantify the computational gains while ensuring that the spatial partitioning used by FAISS did not degrade the integrity of the proposed ADFs.
+As summarized in Table 4, the implementation of FAISS (using an IVF-Flat index with nprobe=16) yielded results identical to the Brute Force method (under tolerance threshold of 200 meters) across all primary performance metrics: Precision (82.94%), Recall (62.15%), and F1-score (71.05%).
+This demonstrates that the approximation error of the IVF index is negligible under a threshold of 200 m in this case, successfully capturing the relevant neighbor set ($k=100$) required for accurate density estimation.
+The most significant finding lies in the computational throughput.
+The Brute Force approach required approximately 10.82 ms per query point, whereas the FAISS-optimized system reduced this to 0.11 ms.
+This represents a nearly 100-fold increase in processing speed.
+Such a reduction is critical for real-time aviation applications, transforming a process that would take minutes into one that executes in seconds, thereby enabling the scalable analysis of large-scale flight datasets.
+Table 4: Comparison table: ADF with FAISS vs. without FAISS
+Precision Recall F1 Average time/query
+with FAISS $82.94\%$ $62.15\%$ $71.05\%$ $0.1145$ ms/query
+without FAISS $82.94\%$ $62.15\%$ $71.05\%$ $10.8184$ ms/query
+4.3 On nprobe and k
+In this section, we evaluate the impact of the index search depth, defined by the nprobe parameter, on both the classification performance and computational efficiency of the Adaptive Density Function (ADF).
+The parameter $k$ was fixed at 100 to ensure a sufficiently large neighborhood for the local density kernels, while nprobe was varied from 4 to 256 to observe the approximation error introduced by the Inverted File (IVF) structure.
+As demonstrated in Table 5, the classification metrics (Precision, Recall, and F1-score) remain remarkably stable across different nprobe values.
+Even at a low search depth of nprobe=8, the system achieves an F1-score of 71.05%, identical to the score at the highest search depth of 256.
+This suggests that the high-density regions identifying Points of Interest (POIs) are sufficiently distinct that even a coarse approximate search effectively captures the primary density contributors.
+However, the computational cost scales significantly with the search depth.
+While nprobe=4 provides a latency of approximately 0.1255 ms per query, increasing the depth to 256 raises the latency to 0.9571 ms, a nearly eight-fold increase in processing time for negligible gains in classification accuracy.
+It is worth noting that at very low nprobe values, diminishing returns in speed were observed, likely due to fixed computational overheads in the Python-C++ interface.
+Based on these results, nprobe=16 was selected as the optimal operating point for the remainder of the study, providing the best compromise between mathematical fidelity and real-time processing requirements.
+Table 5: Comparison table: nprobe selection
+nprobe Precision Recall F1 Average time per query
+4 $82.92\%$ $62.14\%$ $71.04\%$ $0.1255$ ms/query
+8 $82.94\%$ $62.15\%$ $71.05\%$ $0.1484$ ms/query
+16 $82.94\%$ $62.15\%$ $71.05\%$ $0.1596$ ms/query
+64 $82.94\%$ $62.15\%$ $71.05\%$ $0.2920$ ms/query
+256 $82.94\%$ $62.15\%$ $71.05\%$ $0.9571$ ms/query
+Note: The threshold is set to 200 meters, k=100.
+Following the index optimization, the neighborhood size $k$ was evaluated to determine its influence on the ADF smoothing effect and classification accuracy.
+As shown in Table 6, the $k$ parameter was varied from 50 to 150 with nprobe fixed at 16.
+Similar to the search depth findings, the system displays high resilience to changes in $k$, with the F1-score fluctuating by less than 0.03%.
+This stability indicates that the regional density peaks are sufficiently dense that the core POI classification is not highly sensitive to the peripheral neighbors included in the kernel summation.
+A minor increase in precision and F1-score was observed when moving from $k=50$ to $k=150$, while precision remained remarkably stable (fluctuating by less than 0.03%).
+suggesting that a larger neighborhood helps bridge small gaps in the density field, resulting in more cohesive POI extraction.
+However, increasing $k$ beyond 100 yielded diminishing returns, with a linear increase in computational latency from 0.14 ms to 0.17 ms per query.
+This latency increase is attributed to the higher dimensionality of the distance and score vectors processed during the kernel computation phase.
+Consequently, $k=100$ was finalized as the optimal neighborhood size based on the speed and F1-score as it maximizes recall without incurring the over-smoothing or computational costs associated with higher $k$ values.
+Table 6: Comparison table: k value selection
+k Precision Recall F1 Average time per query
+50 $82.97\%$ $62.10\%$ $71.04\%$ $0.1435$ ms/query
+100 $82.94\%$ $62.15\%$ $71.05\%$ $0.1596$ ms/query
+150 $82.95\%$ $62.16\%$ $71.07\%$ $0.1724$ ms/query
+Note: The threshold is set to 200 meters, nprobe=16.
+Ultimately, the combination of nprobe=16 and $k=100$ establishes a high-performance configuration capable of processing trajectory points at sub-millisecond speeds while maintaining the mathematical fidelity of a brute-force approach.
+5 Comparative Evaluation and Related Works
+5.1 A Simple Comparison with Classical k-Nearest-Neighbors method(KNN)
+For the benchmark comparison study with existing traditional methods, we compared the proposed framework with the standard K-Nearest-Neighbors (KNN) as a canonical baseline.
+The K-Nearest-Neighbors (KNN) baseline was implemented as an unsupervised density estimator operating in three-dimensional Earth-Centered, Earth-Fixed (ECEF) coordinates.
+This method serves as a purely geometric proximity baseline, where the local density at any trajectory point is inversely proportional to the mean distance of its closest neighbors in the reference POI dataset.
+Unlike supervised classification, this approach identifies Points of Interest (POIs) by selecting trajectory points that exhibit the highest local density (lowest mean distance) relative to the global distribution of the flight.
+As implemented in the experimental pipeline, flight coordinates are first transformed to Cartesian ECEF space to ensure consistent Euclidean distance metrics across varying flight levels.
+For each query point, the algorithm identifies the set of neighbors and computes the density proxy as:
+$D(p)=\frac{1}{k}\sum_{q\in\mathcal{N}_{k}(p)}|p-q|_{2}$
+where lower values of $D(p)$ correspond to higher local density.
+A binary POI mask is then generated by applying a threshold at the 75th percentile of the observed distances (selecting the 25% of points with the highest proximity).
+The performance of this baseline was evaluated across a sweep of values to assess sensitivity to the neighborhood scale.
+Table 7: KNN Baseline Performance Metrics (3D ECEF, 200m Threshold)
+Parameter k Precision Recall F1-Score Latency (ms/query)
+25 53.91% 95.40% 68.89% 0.3671
+50 53.70% 95.09% 68.64% 0.3361
+75 53.46% 94.81% 68.37% 0.3499
+500 51.31% 92.22% 65.93% 0.4753
+The quantitative results in Table 7 highlight a significant trade-off in the KNN approach.
+While the algorithm achieves an exceptionally high recall (exceeding 95% at $k=25$), it suffers from poor precision ($\approx 50\sim 55\%$).
+This behavior results in a large number of False Positives (exceeding 28,000 points per run), suggesting that while geometric proximity is a necessary condition for identifying POIs, it is not a sufficient one; the lack of score-modulation in KNN leads to the inclusion of high-density but low-significance background points.
+Furthermore, the increase in k to 500 leads to a degradation in both precision and recall, confirming that over-extending the neighborhood window introduces excessive noise into the local density estimation.
+5.2 Relation to Kernel Density Estimation
+Classical KDE [30] evaluates a global sum over all points using fixed or locally adaptive bandwidths.
+In contrast, the Adaptive Density Field (ADF) is a query‑conditioned, local aggregation operator: for each query location ADF selects a bounded neighbor set (via k‑NN retrieval) and aggregates score‑modulated kernels centered on those neighbors.
+Although ADF shares mathematical components with adaptive KDE [31] (both use variable bandwidths and local weighting), their objectives differ: adaptive KDE methods estimate probability densities from samples and typically rely on all observations, whereas ADF constructs an application‑driven influence field.
+In ADF, bandwidths are modulated by externally supplied influence scores (or learned mappings) and the kernel support is defined by the query‑dependent neighbor set, so the operator preserves spatial structure and application semantics rather than producing a normalized probability density.
+Consequently, ADF does not estimate a traditional probability density, but constructs a continuous influence field that preserves spatial structure and application-specific relational information.
+The table shown below summarizes the direct comparison between variable‑bandwidth KDE and the proposed ADF.
+Table 8: Conceptual and Structural Comparison between VBKDE and ADF
+Feature Traditional VBKDE Proposed ADF (Attention-Based)
+Input Data Points only $P$ Points $P$ with application-specific scores; extensible to geometry-conditioned relations $G$
+Bandwidth $\sigma$ Function of local point density $\rho(P)$ Function of the interaction between $G$ and $P$
+Kernel Shape Almost always Isotropic Isotropic in this study; anisotropic supported by framework
+Logic "How many points are near this spot?" "How much does this POI matter to this specific flow?"
+ANN-accelerated KDE method [37] primarily focus on speeding up density estimation.
+In contrast, ADF integrates approximation directly into the operator: k‑NN selection serves as top‑k attention pruning rather than a mere post hoc speedup.
+In this way, ADF reframes adaptive kernels as a geometry‑preserving attention operator, where sparsification is intrinsic and kernel choices and score-bandwidth mappings remain design degrees of freedom.
+5.3 Relation to the Attention Mechanism
+The proposed ADF method was inspired by the Attention Mechanism [35].
 While attention mechanisms are typically defined over discrete tokens in a learned latent space, ADF operates directly in continuous metric space.
@@
 As a result, ADF should be viewed not as a rebranding of Transformer attention, but as a geometric generalization of attention to continuous spatial domains.
-Table 1
-Here’s the direct correspondence:
+Table 9: Structural Mapping: ADF Components to Attention Mechanism
 ADF Component Attention Component Meaning
@@
 POI scores $s_{i}$ Value vectors $V_{i}$ “What do we aggregate?”
-Gaussian kernel (Unormalized) attention weights “How strongly do we attend to each neighbor?”
-FAISS k‑NN Top‑k attention pruning “Only attend to the nearest keys.”
-This correspondence is structural rather than metaphorical:
-ADF satisfies the abstract definition of attention as a metric-induced aggregation mechanism over continuous space.
+Gaussian kernel Attention weights “How strongly do we attend to each neighbor?”
+FAISS k-NN Top-k attention pruning “Only attend to the nearest keys.”
+This correspondence is structural rather than metaphorical: ADF satisfies the abstract definition of attention as a metric-induced aggregation mechanism over continuous space.
 The attention analogy is used as a structural lens rather than a claim of architectural novelty.
-Unlike Transformer’s attention, which operates in a learned latent space, ADF attention is geometry-preserving and grounded in physical distance.
-No projection matrices are required; relevance emerges directly from the spatial metric.
-Moreover, ADF attention is metric-induced and geometry-preserving, requiring no learned projection matrices.
-2.6.3 Why This Matters
-The suggested ADF is a geometric attention:
-  • Query = location where you want influence
-  • Keys = POI positions
-  • Values = POI scores
-  • Kernel = attention weights
-  • FAISS = top‑k attention pruning
-This built a spatial attention mechanism.
+5.4 Other Related Works
+Pattern matching and spatial query systems return discrete matches under complex constraints and rely on different indexing strategies (e.g., IR‑trees) [13], distinguishing them from continuous influence operators like ADF.
+Control‑based adaptive retrieval (e.g., PIDKNN) adapts search radius per query and offers an alternative to ANN retrieval [26].
+Density‑based clustering methods such as OPTICS reveal hierarchical density structure and avoid a single global density threshold, but they operate in batch and output cluster labels rather than continuous, query‑conditioned influence fields [12].
+Adaptive KDE and k‑NN density estimators form the statistical backbone for local bandwidth selection and smoothing [27].
+Hybrid kNN–kernel methods have been applied to spatio‑temporal clustering and activity detection [24, 19], but they operate as dataset‑level estimators rather than per‑query operators.
+Query‑conditioned influence algorithms and scalable spatial query systems address per‑query efficiency [25, 2], yet they typically define influence via density contribution or discrete matching rather than as a score‑modulated attention operator.
+ANN‑accelerated KDE pipelines demonstrate practical scalability on large datasets [32], but in those works ANN is an implementation detail rather than an intrinsic component of the operator.
+ADF differs by (i) treating neighbor selection and sparsification as part of the operator, (ii) modulating kernel bandwidths with external scores, and (iii) framing the aggregation as geometry‑preserving attention.
+By embedding ANN-based sparsification into the operator itself, ADF addresses the computational bottleneck of high-resolution density modeling in a way that remains theoretically consistent with geographic principles of locality.
+5.5 Why This Matters
+The suggested ADF is best understood as a geometric attention mechanism: a query‑conditioned, metric‑driven aggregation over continuous space.
+In this sense, ADF implements a spatial attention operator rather than a traditional kernel estimator.
 While ADF shares mathematical components with adaptive KDE and ANN-accelerated density estimation, its contribution is not a new kernel estimator per se.
 Instead, ADF explicitly formulates these components as a geometric attention operator, where approximation, sparsification, and metric structure are intrinsic to the definition rather than implementation details.
-Importantly, the attention analogy is used as a structural lens rather than a claim of architectural novelty.
-ADF does not seek to replace or subsume Transformer-style attention, but to extend the abstract notion of query-conditioned weighted aggregation to continuous metric spaces.
 Overall, ADF should be understood as a framework-level operator rather than a fixed algorithm, with emphasis on structure, scalability, and geometric grounding.
-3. Implementation and results
-Given a large set of aircraft trajectories, local motion irregularities are first detected as Points of interest (POIs).
-These POIs are sparse, noisy, and heterogeneous.
-We aim to aggregate them into a continuous spatial field to reveal Zones of Influence (ZOIs): regions where systematic maneuvering, airspace structure, or control-induced behavior recurrently concentrates.
-In many spatial decision problems (e.g., flight routing, navigation, risk exposure), the agent does not respond to absolute field intensity, but to relative dominance along its trajectory.
-Zones of Influence (ZOIs) are therefore introduced as trajectory-conditioned dominance regions, representing spatial areas where the accumulated influence field locally dominates competing influences relative to the agent’s path.
-3.1. Study area and data description
-We used a nationwide flight trajectory dataset over mainland China for two multi-periods: 2024-11-10 to 2024-11-22 and 2024-12-05 to 2024-12-15 to extract the POIs.
-There are a total of 1.8 millions of POIs extracted from the dataset using the motion-based procedure described in Section 3.2.
-For the case study on trajectory-conditioned ZOIs, we selected a single day and airport: the Chengdu Shuangliu airport (CTU) on 2024-12-16.
-All trajectories intersecting the Chengdu region on that day were used to construct a local example of the Adaptive Density Field (ADF) and to illustrate ZOI extraction along individual flights.
-The ADF framework is not restricted to flight trajectories and can be instantiated in any domain where localized motion irregularities or events are observed.
-Likewise, the POI definition is not universal; the construction adopted here is a concrete example tailored to aircraft motion.
-3.2. POI construction
-To collect the POI data, we applied a mixture of motion-based POI detection methods, detailed in the Appendix.
-However, POIs are derived from trajectory segments where a physics‑based motion model fails to predict future positions accurately.
-The processing pipeline operates at the individual flight level.
-For each flight, raw WGS‑84 positions (longitude, latitude, altitude) are first converted to a local East–North–Up (ENU) coordinate system.
-The trajectory is then represented as a dictionary keyed by flight identifier, with three aligned time series:
-  • coords: longitude, latitude, altitude
-  • vel: velocity components (velocity on each axis)
-  • dt: timestamp interval
-To estimate future aircraft positions, I applied a physics-based interpolation model that blends two motion predictors:
+6 Discussion
+The proposed Adaptive Density Field (ADF) framework provides a flexible, scalable mechanism for aggregating sparse, heterogeneous points of interest (POIs) into a continuous spatial field.
+By coupling trajectory-conditioned analysis with POI extraction as a downstream instantiation of ADF, this methodology captures both globally recurrent patterns and locally significant deviations.
+6.1 Future application and extensions
   1. 1.
-    Constant-Acceleration (CA) model: reliable for nearly straight trajectories
-  2. 2.
-    Cubic Hermite Spline interpolation: smooth and accurate for curved motion
-the blending weight is determined by the local curvature of the trajectory [lee2007trajectory]: Low curvature indicates motion is nearly straight, hence CA weights more, and vice versa.
-This adaptive combination produces a more stable and realistic prediction thanm using either method alone.
-To evaluate the quality of the predicted positions, I compute a time‑normalized Mahalanobis loss (This metric captures both the magnitude and directional structure of prediction errors.) for each flight.
-This metric captures not only the magnitude of prediction errors but also their directional structure, covariance, and temporal spacing.
-After computing the time‑normalized Mahalanobis loss for each flight, the next step is to identify Points of Interest (POIs)—locations where the prediction error is unusually high [chandola2009anomaly].
-These points often correspond to sharp maneuvers, abnormal motion, or sensor irregularities, and they serve as valuable markers for downstream analysis.
-However, in this case, a POI does not necessarily correspond to an actual infrastructure feature; it simply marks a point where the motion deviates significantly from our prediction.
-It can be regard as an example tailored POI definition under certain circumstances.
-Later we normalize the loss scores and filter the POI with a threshold and exported the data into csv format.
-3.3. ADF construction
-We instantiated the ADF framework on the nationwide POI dataset by specifying the coordinate system, kernel parameterizations, and neighbor search configurations.
-All POI locations were converted from WGS84 geodetic coordinates to Earth-Centered, Earth-Fixed (ECEF) Cartesian coordinates to ensure metric consistency.
-POI scores were treated as non-negative scalar probability attention values and were used directly without additionaly normalization.
-For the scalable neighbor retrieval, a FAISS inverted file (IVF) index was introduced and constructed over the ECEF coordinates.
-The index was trained with 4096 clusters ($2^{12}$) clusters, and queries probed the 16 nearest clusters.
-For each query point the top $k=100$ approximate nearest neighbors were retrieved.
-The global spatial scale parameter was fixed to $\sigma_{0}=500$ meters for all experiments.
-ADF values were evaluated using the formulation defined in Section 2, without further modification.
-Unless otherwise stated, these parameter choices were held constant throughout the case study.
-3.4. Field evaluation
-To analyze the ADF along individual trajectories in Chengdu, we:
-  1. 1.
-    Convert the coordinates: transform all points from geodetic to ECEF.
-  2. 2.
-    Evaluate ADF: compute $ADF(x_{t})$ at each trajectory point $x_{t}$ using the FAISS-accelerated framework.
-  3. 3.
-    Re-attach geodetic coordinates: join ADF values back to $(lon,lat,alt)$ for downstream analysis.
-This allows us to examine how aircraft trajectories intersect regions of concentrated POIs.
-3.5. ZOI extraction
-Given the ADF evaluated along a trajectory, we extract Zones of Influence (ZOIs) as segments where the field intensity dominates a baseline [zheng2015trajectory] defined by the flight’s own experience.
-Let $\{F_{t}\}$ denote the sequence of ADF values along a single trajectory.
-We define a robust baseline level as the trajectory‑wise median [rousseeuw1993alternatives]:
-$F_{base}={median}_{t}(F_{t})$
-A point $t$ is labeled as belonging to a ZOI if
-$F_{t}\geq\alpha F_{base}$
-where $\alpha>1$ is a dominance factor.
-In the Chengdu case study, we set $\alpha=1.3$, chosen empirically to balance sensitivity and specificity in detecting high-density segments along individual trajectories.
-The resulting binary mask partitions the trajectory into ZOI and non‑ZOI segments.
-Trajectories primarily in low-density regions may still exhibit short ZOIs, while trajectories in busy airspace require higher ADF values to exceed the baseline.
-Figure 2: Trajectory ADF analysis and ZOI extraction
-Visualization
-To illustrate the ADF and ZOI extraction:
-  • Trajectories are rendered as white lines, tracing the actual flight paths through the airspace.
-  • Red dots mark the extracted Zones of Influence (ZOIs)—trajectory points where the ADF value exceeds the dominance threshold relative to the flight’s own baseline.
-  • The remaining trajectory points are colored by their Adaptive Density Field (ADF) values (within range 0~30), using a continuous colormap from blue (low intensity) to red (high intensity), with values ranging from 0 to 80.
-  • Layered maps allow toggling between POIs, ADF, ZOIs, and trajectory paths, enabling flexible inspection of each component.
-These visualizations demonstrate the relationship between the continuous ADF field and trajectory-conditioned ZOIs.
-High-ADF segments generally coincide with operationally significant areas such as approach corridors, holding patterns, and vectoring zones in the Chengdu terminal.
-ZOI points, highlighted in red, mark locations where the trajectory locally exceeds its baseline field intensity; some of these points overlap dense ADF regions, representing recurrent maneuvering, while others appear in otherwise sparse areas, potentially reflecting transient anomalies, unique pilot actions, or data artifacts.
-By layering ADF intensity, ZOI labels, and raw POIs, the visualization distinguishes between structurally recurrent behavior and isolated events, offering a nuanced perspective on spatial influence and motion irregularities.
-4. Discussion
-The proposed Adaptive Density Field (ADF) framework provides a flexible, scalable mechanism for aggregating sparse, heterogeneous points of interest (POIs) into a continuous spatial field.
-By coupling trajectory-conditioned analysis with ZOI extraction as a downstream instantiation of ADF, this methodology captures both globally recurrent patterns and locally significant deviations.
-4.1. Future application and extensions
-  1. 1.
-    Operational Airspace Management: The ADF-ZOI pipeline we presented in the example implementation section could assist air traffic controllers in identifying target regions depending on the specific POI definition, such as regions of frequent maneuvering or congestion, enabling data-driven decisions for route planning, holding pattern optimization, and as well as safety monitoring.
+    Operational Airspace Management: The ADF-POI pipeline we presented in the Section 3 could assist air traffic controllers in identifying target regions depending on the specific POI definition, such as regions of frequent maneuvering or congestion, enabling data-driven decisions for route planning, holding pattern optimization, and and safety monitoring.
   2. 2.
     Predictive Trajectory Analysis: Beyond post-hoc evaluation, the framework can be integrated with predictive models for trajectory planning, anomaly detection, or risk exposure assessment.
-By defining the POIs, it is possible to derive trajectory-conditioned ZOIs to provide adaptive thresholds for identifying unusual events relative to expected motion patterns.
+By defining the POIs, it is possible to derive trajectory-conditioned POIs to provide adaptive thresholds for identifying unusual events relative to expected motion patterns.
   3. 3.
-    Cross-Domain Applicability: While this study focuses on aircraft trajectories due to the initial problem setting, the methodolody is domain-agnostic.
-Any scenario involving spatiotemporal events, as long as the spatial characteristics are stable and fixed (in other words, not dynamic), such as maritime traffic, pedestrian movement, vehicle flows, or wildlife tracking – we expect ADF-based spatial attention is able to identify zones of concentrated activities.
+    Cross-Domain Applicability: While this study focuses on aircraft trajectories due to the initial problem setting, the methodology is domain-agnostic.
+Any scenario involving discrete spatiotemporal observations with underlying metric regularities, as long as the spatial characteristics are stable and fixed (in other words, not dynamic), such as maritime traffic, pedestrian movement, vehicle flows, or wildlife tracking – we expect ADF-based spatial attention can identify zones of concentrated activities.
   4. 4.
     Integration with Semantic Context: Future work could embed semantic labels, such as airspace type, weather conditions, or operational procedures, into the ADF formulation, yielding geo-semantic density fields.
-This way, interestingly, aligns even more closely with the application of attention mechanism in spatiotemporal analysis.
+This way, interestingly, aligns even more closely with the application of the attention mechanism in spatiotemporal analysis.
   5. 5.
-    Real-Time Applications: With optimizations to indexing, neighbor search, or GPU-accelerated computation, the ADF framework could support near-real-time monitoring of live trajectories in drones, robotics, and other autonomous systems, allowing dynamic ZOI identification and interactive visualization for operational decision-making.
-4.2. Methodological Reflections
+    Real-Time Applications: With optimizations to indexing, neighbor search, or GPU-accelerated computation, the ADF framework could support near-real-time monitoring of live trajectories in drones, robotics, and other autonomous systems, allowing dynamic POI identification and interactive visualization for operational decision-making.
+6.2 Methodological Reflections
   • The trajectory-conditioned approach ensures that high-density areas are interpreted relative to the agent’s path, avoiding misleading conclusions based solely on absolute field intensity.
-  • Parameter choices, such as kernel bandwidth and dominance factor, influence sensitivity and specificity; future studies could explore adaptive or data-driven tuning to optimize ZOI detection across heterogeneous datasets.
-  • Visualization remains crucial for interpretation: layered maps combining POIs, ADF intensity, and ZOI labels provide immediate insight into both recurrent behavior and outlier events, but higher-dimensional visualizations or interactive dashboards could further enhance understanding.
+  • Parameter choices, such as kernel bandwidth and dominance factor, influence sensitivity and specificity; future studies could explore adaptive or data-driven tuning to optimize POI detection across heterogeneous datasets.
+  • Visualization remains crucial for interpretation: layered maps combining POIs, ADF intensity, and POI labels provide immediate insight into both recurrent behavior and outlier events, but higher-dimensional visualizations or interactive dashboards could further enhance understanding.
   • This study represents a conceptual framework, and some details may therefore be approximate, incomplete, or provisional.
-5. Conclusion
+7 Conclusion
 This work introduces Adaptive Density Fields (ADF) as a formulation-level geometric attention operator for scalable aggregation of sparse spatial events in continuous metric space.
@@
 To demonstrate the utility of this formulation, we presented an example instantiation in the context of aircraft trajectory analysis.
-Motion-derived Points of Interest were aggregated into a continuous ADF, and trajectory-conditioned Zones of Influence (ZOIs) were extracted using a relative dominance criterion along individual paths.
-The Chengdu case study illustrates how the resulting fields reveal recurrent airspace structures while distinguishing localized, trajectory-specific deviations.
+Motion-derived Points of Interest were aggregated into a continuous ADF, and trajectory-conditioned Points of Interest (POIs) were extracted using a relative dominance criterion along individual paths.
+The case study illustrates how the resulting fields reveal recurrent airspace structures while distinguishing localized, trajectory-specific deviations.
 Beyond aviation, the ADF framework is domain-agnostic and applicable to a wide range of spatial and spatiotemporal settings, including urban mobility, maritime traffic, robotics, and environmental monitoring.
@@
 Overall, ADF provides a robust and extensible foundation for scalable, interpretable spatial attention, bridging geometric structure and efficient computation in complex spatial environments.
+Notes
+  1. 1.
+    The raw dataset used in the experiments is proprietary and cannot be made public.
+  2. 2.
+    All code is publicly available at: https://github.com/JaimeFine/adaptive-density-field.
+Acknowledgements
+The author would like to thank Dr.
+Yunxiang Han (Sichuan University) for valuable guidance and constructive feedback during the development of this work.
+This research also made use of open-source scientific software, including Python and R for data processing and analysis, and the Leaflet library with OpenStreetMap basemaps (© OpenStreetMap contributors) for interactive geospatial visualization.
+Figures and manuscript preparation were conducted using .
+Generative artificial intelligence
+This manuscript benefited from generative AI tools in limited, non-substantive ways.
+ChatGPT (version 4o-mini) and Gemini 3 were used to improve language clarity and fluency.
+Microsoft Copilot (search mode) was employed to assist with citation discovery and verification.
+All conceptual content, analysis, and argumentation were developed by the author.
 Appendix: Physics-informed Trajectory POI Detection Pipeline
 The appendix provides implementation-level details and mathematical formulations supporting the methods described in the main text. Citations follow the same reference list as the main body.
-1 Preprocessing the Flight Data
-1.1 Coordinate Conversion [hofmann2012global]: WGS84 Geodetic to ECEF
+1. Preprocessing the Flight Data
+1.1. Coordinate Conversion [16]: WGS84 Geodetic to ECEF
 Given:
@@
 Hence we get the ENU coordinates.
-1.2 Coordinate Conversion: ECEF to ENU Conversion
+1.2. Coordinate Conversion: ECEF to ENU Conversion
 Pick a reference point (the origin of the local ENU frame in this case) with geodetic coordinates $(\varphi_{0},\lambda_{0},h_{0})$, and compute its ECEF coordinates $(x_{0},y_{0},z_{0})$ using the same equations as above.
@@
 This is the standard ECEF → ENU transformation used in geodesy and navigation.
-1.3 Creating a Dictionary
+1.3. Creating a Dictionary
 To organize per‑flight data extracted from each GeoJSON file, we build a dictionary where each flight ID maps to three lists:
@@
 In practice, we use a defaultdict so each new flight_id automatically initializes this structure.
-2 Position Prediction
+2. Position Prediction
 To estimate future aircraft positions, I applied a physics‑based interpolation model that blends two motion predictors:
   1. 1.
-    Constant‑Acceleration (CA) model [bar2001estimation] — reliable for nearly straight trajectories
+    Constant‑Acceleration (CA) model [3] — reliable for nearly straight trajectories
   2. 2.
-    Cubic Hermite Spline interpolation [de1978practical] — smooth and accurate for curved motion
-The spline utilizes local velocity vectors as tangents at each waypoint, providing a geometrically consistent path [virtanen2020scipy] that complements the CA model’s acceleration-based predictions."
+    Cubic Hermite Spline interpolation [11] — smooth and accurate for curved motion
+The spline utilizes local velocity vectors as tangents at each waypoint, providing a geometrically consistent path [36] that complements the CA model’s acceleration-based predictions."
 The blending weight is determined by the local curvature of the trajectory: - Low curvature → motion is nearly straight → CA dominates
@@
 This adaptive combination produces a more stable and realistic prediction than using either method alone.
-2.1 General Prediction
+2.1. General Prediction
 For each flight:
@@
 This yields a smooth, curvature‑aware prediction for each flight.
-2.2 Loss Computation
-To evaluate the quality of the predicted positions, I compute a time‑normalized Mahalanobis loss for each flight.
+2.2. Loss Computation
+To evaluate the quality of the predicted positions, I compute a time‑normalized Mahalanobis loss [22] for each flight.
 This metric captures not only the magnitude of prediction errors but also their directional structure, covariance, and temporal spacing.
 The loss is computed in four main steps:
-1. Extract Prediction Residuals
+Extract Prediction Residuals
 For each flight, I compare the predicted positions $\hat{p}_{i}$ with the actual converted coordinates $p_{i}$:
@@
 This removes global bias and ensures the covariance reflects shape rather than offset.
-2. Estimate Residual Covariance
+Estimate Residual Covariance
 The covariance of the centered residuals is computed as:
 $\Sigma=\operatorname{Cov}(\tilde{r})+\lambda I$
-A small Tikhonov regularization [tikhonov1963solution] term $\lambda=10^{-5}$ stabilizes the inversion of $\Sigma$, especially for nearly collinear trajectories.
+A small Tikhonov regularization [34] term $\lambda=10^{-5}$ stabilizes the inversion of $\Sigma$, especially for nearly collinear trajectories.
 The inverse covariance $\Sigma^{-1}$ defines the Mahalanobis geometry of the error space.
-3. Compute Mahalanobis Distance
+Compute Mahalanobis Distance
 For each residual vector:
@@
 This distance penalizes errors more strongly along directions where the model is normally precise, and less along directions with naturally higher variance.
-4. Normalize by Temporal Spacing
+Normalize by Temporal Spacing
 Because timestamps are not uniformly spaced, each error is scaled by a time‑dependent factor:
@@
 This ensures that predictions made over longer time intervals are not unfairly penalized compared to short‑interval predictions.
-3 POI Detection
+3. POI Detection
 After computing the time‑normalized Mahalanobis loss for each flight, the next step is to identify Points of Interest (POIs)—locations where the prediction error is unusually high.
@@
 The POI detection pipeline consists of three main stages:
-3.1 Normalize the Loss Scores
+3.1. Normalize the Loss Scores
 For each flight, the Mahalanobis losses are rescaled to the interval [0, 1]:
@@
 This normalization ensures that POI detection is relative to each flight’s own dynamics, making the method robust to differences in scale, speed, or noise across flights.
-3.2 Thresholding
+3.2. Thresholding
 Here, I introduced an element called POI score, which indicates how anomalous each point is relative to the rest of the flight.
@@
 It can be adjusted depending on the desired sensitivity of the detection process.
-3.3 Export POIs to CSV
+3.3. Export POIs to CSV
 Each detected POI is stored with: